Thomas Stieltjes Institute for Mathematics

Annual Report 1997

Introduction  1
About the Stieltjes Institute  3
Research Programmes  5
International and National Programmes  15
Ph.D. Programmes  21
Ph.D. Theses  26
Number of Researchers  30
Publications  32
1.1. Number Theory  32
1.2. Algebraic and Analytic Geometry  34
1.3. Topology and Dynamical Systems  38
2.1. Operator Theory and Applications  42

2.2. Lie Groups, Special Functions, Approximation Theory and Complex Analysis

45

2.3. Non-linear Partial Differential Equations  50
2.4. Numerical Analysis  54
3.1. Probability  61
3.2. Statistics  67
4.1. Optimization Theory  75
4.2. Stochastic Operations Research  81
Workshops, Seminars, a.o.  86
Foreign Visitors  94
Research Staff  104
Organization of the Thomas Stieltjes Institute  112

Thomas Jan Stieltjes 1856 – 1894

Thomas Stieltjes Institute for Mathematics

Participants:

– University of Amsterdam

– Free University Amsterdam

– Delft University of Technology

– Eindhoven University of Technology

– University of Leiden

– Erasmus University Rotterdam

Managing Board:

Prof.dr. W.R. van Zwet (RUL), scientific director Drs. F. Bakker (RUL), secretary

Address:

Thomas Stieltjes Institute for Mathematics c/o Mathematical Institute, University of Leiden Niels Bohrweg 1, 2333 CA Leiden, The Netherlands Phone: +31 71 527 7045 / 7042 Fax: +31 71 527 6985 email: Stieltjes@wi.leidenuniv.nl URL: http://www.wi.leidenuniv.nl/math/Stieltjes/

The following abbreviations are used:

CWI Centre for Mathematics and Computer Science, Amsterdam EUR Erasmus University Rotterdam IMAU Institute for Marine and Atmospheric Research, Utrecht KNAW Royal Netherlands Academy of Sciences, Amsterdam KNMI Royal Netherlands Meteorological Institute KUB Tilburg University KUN Nijmegen University NWO Netherlands Organization for Scientific Research, The Hague RIVM National Institute of Public Health and the Environment,

Bilthoven

RL Limburg University, Maastricht RUG University of Groningen RUL University of Leiden RUU University of Utrecht TUD Delft University of Technology TUE Eindhoven University of Technology UT Twente University of Technology, Enschede UvA University of Amsterdam VU Free University Amsterdam WAU Wageningen Agricultural University

Contents

Introduction 1

About the Stieltjes Institute 3

Research Programmes 5

International and National Programmes 15

Ph.D. Programme 21

Ph.D. Theses 26

Number of Researchers 30

Publications 32 1.1. Number Theory . . . . . . . . . . . . . . . . . . . . . . . . . . 32 1.2. Algebraic and Analytic Geometry . . . . . . . . . . . . . . . . 34 1.3. Topology and Dynamical Systems . . . . . . . . . . . . . . . . 38 1.4. Theoretical Foundations of Computer Sciences . . . . . . . . . 39 2.1. Operator Theory and Applications . . . . . . . . . . . . . . . 42 2.2. Lie Groups, Special Functions, Approximation Theory and

Complex Analysis . . . . . . . . . . . . . . . . . . . . . . . . 45 2.3. Non-lineair Partial Differential Equations . . . . . . . . . . . . 50 2.4. Numerical Analysis . . . . . . . . . . . . . . . . . . . . . . . . 54 3.1. Probability . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 3.2. Statistics . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 4.1. Optimization Theory . . . . . . . . . . . . . . . . . . . . . . . 75 4.2. Stochastic Operations Research . . . . . . . . . . . . . . . . . 81

Workshops, Seminars a.o. 86 Foreign Visitors 94 Research Staff 104 Organization of the Thomas Stieltjes Institute 112

Introduction

I’m proud to present the annual report of the Thomas Stieltjes Institute for Mathematics for 1997. This annual report constitutes a break in the short tradition of publishing biennial reports for 1993&1994 and 1995&1996. The reason is that we are presently completing our application for renewal of the formal recognition as a research school by the Royal Netherlands Academy of Sciences for which these reports are needed, and we obviously can not complete a report for 1998 yet. In this application to the Academy we show that:

– the Stieltjes Institute has become an excellent platform for conducting joint research and for organizing joint activities such as workshops and seminars;

– through this collaboration a critical mass has been obtained to ensure a significant impact of our research in many areas of mathematics;

– the Stieltjes Institute conducts a rich programme of educational activities for its graduate students. In short, our application will show that the collaboration of the six university departments that form the Stieltjes Institute has greatly increased the visibility of our mathematical work, both in the Netherlands and abroad. Our collaboration with the Centre for Mathematics and Computer Science at Amsterdam has also been fruitful. We shall therefore submit our application to the Academy with confidence.

Jointly with the other mathematical research schools MRI and EIDMA we submitted a proposal for very substantial funding from the ‘diepte strategie’ programme of the Netherlands Organization for Scientific Research (NWO). Although this proposal was not funded, it was positively received. We intend to intensify our collaboration with MRI and EIDMA and submit a similar proposal during the next phase of this programme.

Honorary doctorates were awarded to I. Gohberg by the Technische Universität Darmstadt and to W.R. van Zwet by the Univerzita Karlova at Prague. On the recommendation of the education committee M.H. Wegkamp received the Stieltjes Prize for his PhD thesis entitled ‘Entropy Methods in Statistical Estimation’.

W.R. van Zwet Scientific Director

About the Stieltjes Institute

The Thomas Stieltjes Institute for Mathematics is a Dutch research institute in mathematics, in which participate the University of Amsterdam, the Free University of Amsterdam, the Delft University of Technology, the Eindhoven University of Technology, the University of Leiden and the Erasmus University of Rotterdam. The Institute collaborates with the Centre for Mathematics and Computer Sciences (CWI) in Amsterdam. The University of Leiden is administrator of the Institute.

The Stieltjes Institute was founded the 12th of November 1992 and carries out research in four main areas of fundamental and applied mathematics: Algebra & Geometry, Analysis, Stochastics and Operation Research. The Institute has a research training programme for Ph.D. students, which started September 1, 1993. In June 1994 the Institute received formal recognition as a research school (onderzoekschool) from the Royal Netherlands Academy of Sciences (KNAW).

Mathematics has many different branches and two different aspects: the fundamental and the applied. However, it is a single coherent body of knowledge. This is demonstrated time and again by sometimes unexpected links between the various parts of mathematics and by the possibilities to apply even the most abstract developments. Also, direct contact with fields of application often generate new challenges for mathematical research. For this reason the Stieltjes Institute does not stress the distinction between fundamental and applied mathematics, and encourages research that crosses the dividing lines between the various parts of mathematics. The Stieltjes Institute combines the expertise of three general, two technical and one economics and medical university. This provides an excellent opportunity for interaction between mathematics and its applications in the life sciences, the natural sciences and engineering, computer science and economics.

4

Research Programmes

Algebra and Geometry

1.1 Number Theory

Programme leader: R. Tijdeman

The research in the programme Number Theory ranges from algebraic number theory and group theory to analytic number theory and ergodic theory. A central theme is the design and application of algorithms in number theory. In algebraic number theory the focus is on algebraic units, class groups and Galois module structures. Methods from class field theory, homological algebra and analytic number theory play an important role. Methods from algebraic number theory and numerical mathematics are applied to factorization methods for large integers. Computational methods from algebraic number theory and Diophantine approximation are applied to determine all integer or even rational points on algebraic curves, and, besides ineffective methods, to study diophantine equations and recurrence sequences. Probabilistic and ergodic theoretical aspects of number expansions are investigated, in particular of β-expansions and several different continued fraction expansions. This is interrelated with the study of Sturmian sequences. In group theory the research deals with representation theory of groups, Dade correspondence theory and extension theory of characters. The coherence, size and intensity of the number theory research has been stimulated by the biweekly Number Theory Seminar. In the national SWON-NWO (groot)project Number Theory De Weger, De Smit and Bosma were active at post-doc positions and Roskam as a graduate student. Additional Stieltjes support made it possible to invite some guests. As an example of the research results one highlight. Some years ago, Schlickewei gave an explicit upper bound for the number of solutions of the linear equation a_{1}x_{1} + ... + a_{n}x_{n} = 1 with unknowns x_{1}, ..., x_{n} from a subgroup G of rank r of the multiplicative group of some algebraic number field K. His upper bound depended on n, the rank r of G and the degree d of K. It was expected that the dependence on d is not necessary. Together with Schlickewei and Schmidt, Evertse succeeded in deriving an upper bound depending only on n and r. A major tool was a further improvement of the so-called quantitative Subspace Theorem as obtained by Schlickewei and Evertse.

1.2. Algebraic and Analytic Geometry

Programme leader: G.B.M. van der Geer

Central research themes are: Algebraic Geometry The study of algebraic varieties with the real numbers as base field. Special interest is given to mappings between real algebraic varieties, cycle classes and homology classes in real algebraic geometry, the underlying real structure of complex varieties and the sudy of complete intersections in differential topology and analytic geometry. Algebraic Geometry of Curves and Abelian Varieties The study of moduli spaces of abelian varieties and curves both in characteristic zero and characteristic p. Also the moduli spaces of vector bundles and K3 surfaces are studied. Curves and varieties over finite fields are studied in relation with coding theory. Mathematical Physics The study of mathematical aspects of quantum field theory and string theory. Special emphasis on the relations with algebraic geometry, such as quantum cohomology, mirror symmetry, moduli space of Riemann surfaces. Topological field theory and manifold invariants. Nonperturbative string theory, string duality and extended objects.

1.3. Topology and Dynamical Systems

Programme leaders: J. van Mill, J.M. Aarts

Dutch topology has a long tradition and enjoys an excellent international reputation. The work of the Dutch topologists covers all of General Topology and its connections with other disciplines. There are ongoing collaborations with distinguished researchers from Canada, the Czech Republic, Poland and the United States and the Dutch group plays an important role in the international organization of topological research. In the current project various topological techniques are applied to research topological dynamics (including topological groups) measure theory and geometric topology. These techniques come from dimension theory, infinitedimensional topology, continuum theory, the theory of (ultra)filters and descriptive set theory.

In the coming period the researched will be concentrated on the following topics:

– The relation between compact homogeneous ANR’s and topological manifolds.

– Measures on compact L-spaces (the so-called Kunen spaces).

– Zerodimensional and pseudocompact topological groups.

– The relationship between dimension theory and the preservation of fixed points under extensions of maps (Ljusternik-Schnirelman).

– Application of continuum theory to topological dynamics and compactification theory.

– Dynamical Systems, Holomorphic dynamical systems, and coupled lattice maps.

1.4. Theoretical Foundations of Computer Sciences

Programme leader: G.Rozenberg

Three important areas of research are: The classic theory of formal languages In particular such fields are investigated as: grammars (contextual grammars, splicing); automata and transducers; combinatorial structure of formal languages; models for natural computation; attribute grammars. The theory of graphs and graph transformations In particular the theory of node-rewriting graph grammars, hypergraph grammars, the logical description of graph transformations, and the hierarchical decomposition of graphs using 2-structures. The theory of concurrent processes In particular Petri nets, traces, vector controlled concurrent systems, transition systems, and the application of Petri nets to model object oriented systems.

Analysis

2.1. Operator Theory and Applications

Programme leaders: M.A. Kaashoek, S.M. Verduyn Lunel

This programme deals with operator theory problems arising from concrete classes of integral and differential equations. Both linear and non-linear equations are studied, and the problems may have a finite dimensional or infinite dimensional character. Typical for the project is a strong interaction with complex function theory; operators are studied in terms of analytic functions (symbols, characteristic matrices, determinant functions). The study of non-stationary versions of normed constrained interpolation problems of Nevanlinna-Pick and Sarason type, the development of a qualitative theory of infinite dimensional dynamical systems, the analysis of the asymptotic behaviour of deterministic and non-deterministic systems, and the study of imaginary powers of unbounded operators are some of the main themes.

2.2. Lie Groups, Special Functions, Approximation Theory and Complex Analysis

Programme leaders: G. van Dijk, T.H. Koornwinder

Central research themes are:

(1) analysis on Lie groups, semisimple symmetric spaces and quantum groups (2) special functions associated with root systems and their interpretation on the above-mentioned structures

(3) analysis, asymptotics, approximation theoretic properties and algorithmic aspects of special functions and orthogonal polynomials in one variable (including the cases of orthogonal rational functions and orthogonality in a Sobolev space)

(4) approximation problems with relation to potential theory

(5) analysis in several complex variables

(6) wavelets

(7) analysis aspects of modern mathematical physics This programme unites a number of themes which are mutually connected and have a stimulating influence on each other. Some themes have a quite algebraic setting, while others belong to classical analysis. Theme (2) dealing with Heckman-Opdam hypergeometric functions and q-analogues like Macdonald polynomials is a highlight of the programme. A new line within theme (1) is the study of canonical representations for Hermitian and para-Hermitian symmetric spaces, and its relation with Berezin quantization.

2.3. Non-linear Partial Differential Equations

Programme leader: L.A. Peletier, Ph.P.J.E. Clément

This programme embraces the study of nonlinear variational problems and evolution equations. The emphasis lies on problems arising in the sciences and in engineering. Traditionally, the group has been very active in the study of nonlinear difffusion problems, especially those involving interfaces or other singularities, and often related to specific applications. However, recently research has branched out into the study of higher order model equations such as the Swift-Hohenberg equation, arising in mathematical physics and nonlinear solid mechanics.

Research takes place in a variety of individual and institutional collaborations. We mention here the Centre for Nonlinear Partial Differential Equations Amsterdam-Delft-Leiden, the NWO initiative on Nonlinear Systems, the HCM Network with groups from Bonn, Florence, Madrid, Paris and Oxford, the ongoing collaborations with the universities of Bordeaux, Lyon, Madrid, Nottingham, Pittsburgh and Trieste and the HCM Network ‘Reaction Diffusion Equations’with Madrid, Edinburgh, Toulouse, Trieste, Athens. The Delft part of the programme is concerned with functional analytic aspects of nonlinear partial differential equations of elliptic and parabolic type as well as coupled systems of such equations. Next to those subjects effort has been put in the study of higher order elliptic equations.

2.4. Numerical Analysis

Programme leaders: P.J. van der Houwen, M.N. Spijker

The programme concerns the construction, analysis and implementation of numerical methods for solving mathematical problems. The research ranges from fundamental to more application oriented. Specific topics studied in the programme comprise the numerical solution of ordinary and partial differential equations, approximation theory and implementational aspects of numerical algorithms for modern computer architectures. Although these subareas are different in nature, there is a strong interaction between these fields of research. The topics studied in the programme belong to the main themes within contemporary numerical analysis. Accordingly, the researchers in the programme have many international contacts with numerical analysis research groups abroad. The mathematical tools used in the analysis of the numerical methods are related to many other parts of mathematics (operator theory, complex analysis and the theory of differential equations).

Stochastics

3.1. Probability

Programme leader: M.S. Keane

Research in this programme is carried out in a very broad range of topics, covering most of the aspects of the mathematical discipline of probability theory and its applications to other mathematical areas, other sciences, and industry and technology. The basic themes of the research comprise among others extreme value theory and applications, optimal stopping, infinitely divisible distributions, stochastic recurrence, stochastic geometry, percolation and particle systems, modelling and simulation of geological structures, branching processes, queueing theory and applications, diffusions, ergodic theory, stochastic dynamical systems, fractal geometry and coding, noncommutative probability theory, stochastic inequalities, and simulation.

Probability theory and its applications develops mathematically precise models for quantitatively describing uncertain situations, and applies these models in order to arrive at optimal or nearly optimal decision procedures. Research carried out in this domain in The Netherlands, and in particular in the Stieltjes Institute, is broad-ranging. Inside of the Institute, there are common interests and collaborations with the programmes Statistics, Stochastic Operations Research, Topology and Dynamical Systems and Number Theory.

3.2. Statistics

Programme leaders: W.R. van Zwet, C.A.J. Klaassen

Semi- and nonparametric models, computationally-intensive methods in statistics, stochastic models in the sciences and industrial statistics are the central themes. These include such divers topics as estimation of functions, extreme value theory, bounds in estimation, image analysis, resampling techniques and statistical quality control.

In this research area, developments are very fast stimulated by the new opportunities for applications of statistics due to the increased and still increasing computing power. No doubt these topics will have a dominant position in international statistical research for the years to come. In each of these fields the Stieltjes Institute has research groups with experts of high international reputation. Close cooperation within the Stieltjes Institute stimulates further the quality and efficiency of the research by use of the complementary expertises. International contacts of the various groups are excellent, but have to be consolidated and extended within the framework of the Institute.

Operations Research

4.1. Deterministic Optimization

Programme leader: A. Schrijver

The programme focuses on the design and analysis of mathematical methods and techniques to discrete mathematical and optimization problems, including algorithms for combinatorial, nonlinear, and integer optimization. In particular, topological, geometric, and algebraic methods are investigated (polyhedral, interior-point, eigenvalue, and homotopy methods).

The problems investigated originate from operations research, computer science, and technology. In the cooperation, the University of Amsterdam and the Free University contribute in expertise of discrete mathematics, CWI of combinatorial optimization and algorithmics, and the Delft University of Technology of interior-point methods for linear and integer programming.

4.2. Stochastic Operations Research

Programme leader: A. Hordijk, J. Wessels

The research in this programme is on stochastic networks (sn). Stochastic networks are networks of entities, with particles residing in and moving between these entities according to stochastic processes. A key example is a queueing network, where the entities are service facilities and the particles customers. In the design of computer-, communication-, and manufactoring systems, the most important criterion presently is quality of service, which is expressed in terms of performance and reliability of the systems in relation to their applications. Stochastic networks provide the mathematical models for the description and analysis of these systems. Due to exciting breakthroughs, the theory of sn is going through a period of feverish activity. The programme contains all main research themes on the key elements: modelling, stability, traffic processes, performance analysis, and control.

The already existing cooperation within this programme will be intensified and extended through a SWON-project, in which the research groups of Boxma (CWI), Hordijk (RUL), De Smit (UT), Tijms (VUA) and Wessels (TUE) participate. These groups have partly complementary, partly overlapping expertise in the main mathematical techniques to analyse sn: queueing theory, Markov decision chains, reliability theory, algorithmic probability, and stochastic simulation.

Research results in 1997 The research in Stochastic Operations Research has in 1997 focused on queueing theory and its applications in modern communication networks. Queueing models have been developed, and analysed, for mobile communication systems and for ATM networks. Special attention has been given to the modelling and analysis of novel traffic features that have been revealed by measurements on, a.o., Ethernet and Internet traffic: Such traffic often appears to exhibit features like self-similarity and long-range dependence. It has been shown that those features can be captured by so-called fluid queues fed by on/off sources, if the on- or off-period distribution of at least one source is heavy-tailed. We have analysed the influence of heavy-tailed input distributions on the tail behaviour of key performance measures like buffer content and busy period. Next to fluid queues, we have also studied ordinary queues with heavytailed service time distributions. The main problems under consideration have been (i) extension of the classical heavy-traffic limit theorems to the case of infinite service time variance, and (ii) the influence of the service discipline on the waiting time tail behaviour. The research on multi-server queueing models with per server dependent arrival streams has been continued. Presently the compensation approach is extended to systems with multiple resources, i.e., a job needs not only a server but some additional scarce resource as well. In the STW project on the real-time performance of database systems a lot of progress has been made. The Optimistic Concurrency Control mechanism to support parallellism has been analysed successfully for a wide range of models. The research on traffic shapers for Asynchronous Transfer Mode networks has concentrated on a two-level shaper. In this model the user accumulates credits which allow him to supply temporarily an increased amount of traffic. In the NWO project Neural networks for planning the possible use of neural networks for decision support in planning problems with limited information is studied. The concentration is on rolling horizon models. The joint research with E. Altman and B. Gaujal on: ”Multimodularity and the control of discrete event systems”, received a VanGogh grant. In this research multimodular functions, introduced in 1985 by B. Hajek, have been investigated. They play a key role in the performance analysis of the admission, routing and polling control of stochastic networks. A characterization of multimodularity has been obtained. The asymptotic analysis of sequences of multimodular functions has resulted in structural properties of optimal policies for telecommunication models. In the joint research with A.A. Yushkevich on: ”Blackwell optimality in Markov decision chains with a Borel state space and unbounded rewards” the first results have been obtained. The joint research with V.A. Malyshev on large space-time scaling limits of the random walk has been continued. The so-called Euler limit paths can be used to analyse micro time insensitive performance measures of the random walk, such as (in)stability. Very often these limit paths are deterministic limits, but they can be stochastic as well. On such example studied in 1997 is the transient random walk on the 2dimensional lattice, where the Euler limit paths ”cycle” off to infinity.

Research at CWI in Stochastic Operations Research: queueing theory; Markov decision processes; performance analysis and control of computer and communication networks; reliability and availability of stochastic networks.

International and National Programmes

European Programmes

– ESPRIT Basic Research

GETGRATS: General Theory of Graph Transformation Systems

Cooperation with: University of Antwerp Belgium, Technische Universität Berlin, Germany, Laboratoire Bordelais de Recherche en Informatique, France, Universität Bremen, Germany, Leiden University, The Netherlands, Università di Pisa, Italy, Università di Roma La Sapienza,Italy Coordinator Stieltjes: H.J. Hoogeboom

APPLIGRAPH: Applications of Graph Transformation

Cooperation with: Universität Bremen, Germany, Rheinisch-Westfälische Technische Hochschule, Aachen, Germany, Universitaire Instelling Antwerpen, Belgium, Technische Universität Berlin, Germany, Leiden University, The Netherlands, Katholieke Universiteit Nijmegen, The Netherlands, Università di Pisa, Italy, Università di Roma La Sapienza, Italy Coordinator Stieltjes: G. Rozenberg

– Marine Science and Technology program (MAST)

MMARIE, Modeling of MARIne Ecology Systems

Cooperation with University of Leuven, Belgium Projectleader: P.J. van der Houwen

Neptune

1995 - 1997 Cooperation with Port & Coastal Limited, U.K.; Delft Hydraulics, The Netherlands; Rijkswaterstaat, The Netherlands; GKSS-Forschungszentrum Geesthacht GmbH, Germany; University of Lancaster, U.K.; University of East Anglia, U.K.

– HCM/TMR

AGE, Algebraic Geometry Europe

1992 - 1998 Coordinators Stieltjes: .P. Murre, A.H.J.M. Van de Ven

Singularities and Interfaces in Nonlinear PDE’s

1994 - 1998 Cooperation with: Bonn, Germany, Florence, Italy, Leiden, The Netherlands, Madrid, Spain, Paris, France, and Oxford, U.K. Coordinator Stieltjes: L.A. Peletier

Nonparametric and Semiparametric Statistical Methods

1995 - 1997 Cooperation with: Utrecht University, The Netherlands and 9 foreign institutes Coordinator Stieltjes: J. Einmahl

Discrete Optimisation and Applications

December 1997 - November 2000 Cooperation with: Rheinische Friedrich-Wilhelms-Universität Bonn, Germany, London School of Economics and Political Science, U.K., University of Oxford, Mathematical InstituteU.K., Université Joseph Fourier, Grenoble, France, Università degli Studi di Roma La Sapienza, Italy, Université Pierre et Marie Curie, Paris, France Project leader: A. Schrijver

Reaction-Diffusion Equations

1994 - 1997 Cooperation with: Heriot-Watt University, Edinburgh, U.K., University of Crete, Greece, Technical University of Athens, Greece, University of Madrid, Spain, University of La Laguna, Spain, Università di Trieste, Italy, Université Toulouse, France Coordinator Stieltjes: Ph.P.J.E. Clément

European School of Group Theory

1996 - 1998 Cooperation with France, Italy, Denmark, Spain, Germany Coordinator Stieltjes: E.M. Opdam

Nonlinear Parabolic Partial Differential Equations: Methods and Applications

1998 - 2001 Cooperation with: Bath, U.K., Bonn, Germany, Madrid, Spain, Paris, France, Rome, Italy, and Tel Aviv, Israel Coordinator Stieltjes: J. Hulshof

– INTAS

Hilbert and Krein space operators: extension problems and functional models

June 1996 - October1997 Coordinator Stieltjes: M.A. Kaashoek

Nonparametric statistical inference: asymptotic theory and implementation

June 1996 – June 1998 Cooperation with: Mathematical Research Institute, The Netherlands, Universität Bielefeld, Germany, Belarus Academy of Sciences, Minsk, Belarus, Steklov Mathematical Institute, Moscow, Russia, Institute for Problems of Information Transmission, Moscow, Russia, Mathematical Institute, St. Petersburg, Russia, Razmadze Mathematical Institute, Tbilisi, Georgia. Project leader: W.R. van Zwet

Computer-intensive methods in statistical inference and mathematical finance: asymptotic theory and implementation

December 1995 – June 1997 Cooperation with: Mathematical Research Institute, The Netherlands, Universität Bielefeld, Germany, Belarus Academy of Sciences, Minsk, Belarus, Steklov Mathematical Institute, Moscow, Russia, Institute for Problems of Information Transmission, Moscow, Russia, Mathematical Institute, St. Petersburg, Russia, Razmadze Mathematical Institute, Tbilisi, Georgia, Humboldt Universität Berlin, Germany, Institut für Stochastik, Berlin, Germany, University of Århus, Denmark. Project leader: W.R. van Zwet

Limit theorems for stochastic discrete event systems

October 1994 - October 1997 Cooperation with: Baccelli, INRIA, France, Schassberger, Braunschweig, Germany, Suhov, Cambridge, U.K., Dobrushin, Moscow, Russia, Borovkov, Novosibirsk, Russia Coordinator Stieltjes: O.J. Boxma

– Euclid

CALMA, Solution methods for hard combinatorial optimization problems

Cooperation with: TUE, The Netherlands, Wageningen Agricultural University, The Netherlands, TNO, The Netherlands CERT, Toulouse, France, Kings’ College, London, U.K., University of East Anglia, Norwich, U.K. Project leader: C. Roos

National Programmes

– SWON/NWO (Netherlands Foundation for Mathematics/ Netherlands Organization for Scientific Research)

Computationally Intensive Methods in Stochastics

1993 - 1998 Participants: Thomas Stieltjes Institute for Mathematics, CWI and the Mathematical Research Institute, The Netherlands Coordinators Stieltjes: L.F.M. de Haan, M.S. Keane, C.A.J. Klaassen and W.R. van Zwet

Algebraic curves and Riemann surfaces

1993 - 1998 Participants: UvA, RUU Coordinator Stieltjes: G. van der Geer

Lie theory and special functions

September 1994 - December 1998 Participants: RUL, UvA, RUU, KUN, UT, CWI Coordinators Stieltjes T.H. Koornwinder and E.M. Opdam

Number Theory

1995 - 2000 Participants: RUL, UvA, RUU and University of Groningen Coordinator Stieltjes: R. Tijdeman

Nonlinear Transport Phenomena in Porous Media, theme: Filtration Theory

Participants: CWI, TUD, LUW Coordinator Stieltjes: J. Hulshof

Dynamical systems and pattern formation

September 1997 - August 1998 Participants: VU, RUU, CWI, UU, RUG, TUD Coordinators Stieltjes: L.A. Peletier and S.M. Verduyn Lunel

Stochastic networks

1997 - 2001 Participants: RUL, CWI, UT, VU, TUE Coordinator Stieltjes: A.J. Hordijk

– NCF (National Computer Facility)

Computational Number Theory

1992 - 2001 Participants: RUL, CWI and University of Groningen Coordinators Stieltjes: R. Tijdeman and H.J.J. te Riele

Domain decomposition in transport-chemistry models for shallow seas

March 1997 - February 1998 Projectleader: P.J. van der Houwen

– HPCN (High Performance Computing and Networking in The Netherlands)

3D transport models for shallow seas

Particpants: Delft Hydraulics, TUD, CWI Projectleader: P.J. van der Houwen

– Divers

Applied Mathematics and Computational Methods

1995 - 1999 Cooperation project between the Netherlands and Indonesia Coordinator Stieltjes: R. Helmers

Optimization of the reloading patters of a nuclear reactor

March 1995 - March 1999 A project of the Optimization Group in Delft, in cooperation with the Interfacultair Reactor Instituut Stieltjes coordinator: C. Roos

Ph.D. Programme

Stieltjes courses/educational activities in 1997/1998

Algebra & Geometry

Instructional Conference on Algebraic Independence

R. Tijdeman Luminy, France, September 29 - October 3, 1997

Number Theory Seminar

P. Stevenhagen, R, Rijdeman Various places, 1997/1998

Intercity-Seminar Geometry

G. van der Geer Various places, 1997/1998

Classical Dimension Theory

J. van Mill Amsterdam, Spring 1998

Seminar Topology

J. van Mill Amsterdam, 1997/1998

Theory of 2-structures

G. Rozenberg Leiden, Fall 1997

Formal Languages and Automata 2

J. Engelfriet Leiden, Fall 1997

Theory of Concurrency 2

H.C.M. Kleijn Leiden, Fall 1997

Seminar DNA Computing

H.J. Hoogeboom, G. Rozenberg Leiden, Fall 1997

Analysis

Operator Theory

M.A. Kaashoek, S.M. Verduyn Lunel Amsterdam, Fall 1997

Seminar Analysis and Linear Operators

M.A. Kaashoek Amsterdam, 1997/1998

Computer algebra algorithms for special functions

T.H. Koornwinder Amsterdam, Fall 1997

Advanced Theory of Functions

J.J.O.O. Wiegerinck Amsterdam, Fall 1997

Finite groups of Lie type

H.T. Koelink, E.M. Opdam Leiden, Spring 1998

International Workshop 1997 on Lie Theory

G.F. Helminck Enschede, December 1997

International Summer European School of Group Theory

G. van Dijk Leiden, June/July 1998

Seminar on Lie Groups

E.P. van den Ban, G.J. Heckman, E.M. Opdam, T.H. Koornwinder Utrecht, Spring 1998

Functional Analytic Methods for Partial Differential Equations

Ph.P.J.E. Clément, B. de Pagter Delft, 1997/1998

Modern Computational Fluid Dynamics

B. Koren Leiden, Fall 1997

Defect Correction and Multigrid Methods

P.W. Hemker Amsterdam, Fall 1997

Numerical Analysis and Dynamical Systems

K.J. in ’t Hout Leiden, Fall 1997

Conference Numerical Analysis

J. Kok et al Woudschoten, September 1997

Stochastics

Large Deviations & Time Series Analysis ^{∗}

F. den Hollander, T. Mikosch Utrecht, Fall 1997

Measure theory, Martingales and stochastic integrals

L. de Haan Rotterdam, Fall 1997

Semiparametric models

C.A.J. Klaassen Amsterdam, Spring 1998

Statistical Modelling of Medical Data & Brownian Motion and the Functional CLT ^{∗}

H. van Houwelingen, G. Hooghiemstra Utrecht, Spring 1997

Mathematical Statistics

J.H.J. Einmahl Eindhoven, Fall 1997

∗ are courses offered at a national scale by the Landelijk AIO-netwerk Stochastiek.

Operations Research

Stochastic Operations Research 1 ^{∗}

H.C. Tijms, L.C.M. Kallenberg Utrecht, Fall 1997

Combinatorial Optimization 1 ^{∗}

W. Kern, A.M.H. Gerards Utrecht, Fall 1997

Stochastic Opeartions Research 2 ^{∗}

J. Wessels, I.J.B.F. Adan, O.J. Boxma Utrecht, Spring 1998

Combinatorial Optimization 2 ^{∗}

J.K. Lenstra, C.P.M. van Hoesel, L. Stougie Utrecht, Spring 1998

Constructive Methods for Transient Behaviour of Stochastic Networks and Applications

F.M. Spieksma Leiden, Spring 1998 ∗ are courses offered at a national scale by the Landelijk Netwerk Mathematische Besliskunde.

Ph.D. Theses

C.O.B. de Mirleau Foundations of functional integration and the Schwinger-Dyson equation Thesis advisor: R. Dijkgraaf

January 7, 1997  University of Amsterdam

P.M. de Zeeuw Acceleration of Iterative Methods by Coarse grid Corrections Thesis advisor: P.W. Hemker

January 22, 1997  University of Amsterdam

R.B. Geskus Estimation of Smooth Functionals with Interval Censored Data Thesis advisor: P. Groeneboom

February 11, 1997  Delft University of Technology

E. Amerik On the geometry of smooth hypersurfaces in projective space Thesis advisor: A.J.H.M. van der Ven

April 10, 1997  Leiden University

M.D. Harris Coin Tossing, Random Mirrors and Dependent Percolation: Three Paradigms of Phase Transition Thesis advisor: M.S. Keane

April 28, 1997  Delft University of Technology

M. Elkenbracht-Huizing Factoring integers with the Number Field Sieve

Thesis advisors: R. Tijdeman

May 14, 1997  Leiden University

W.A. van der Veen Parallelism in the numerical solution of ordinary implicit differential equations Thesis advisor: P.J. van der Houwen

May 21, 1997  University of Amsterdam

J. van Hamel Algebraic cycles and topology of real algebraic varieties Thesis advisor: J. Bochnak

May 26, 1997  Free University Amsterdam
C.M. Schuhmann  Morphisms between Fano threefolds

Thesis advisor: A.J.H.M. van der Ven

May 27, 1997  Leiden University

F.A.J. Straetemans Stability and Resolvent Conditions for Discretizations of Initial Value Problems Thesis advisor: M.N. Spijker

May 29, 1997  Leiden University

H. Boender Factoring Large Integers with the Quadratic Sieve Thesis advisor: R. Tijdeman

June 10, 1997  Leiden University
P.G.J. Nijsse  Stable Vector Bundles on Quintic Surfaces

Thesis advisor: A.J.H.M. van der Ven

June 12, 1997  Leiden University

J.F.A.A. Nagel The image of the Abel-Jacobi map for complete intersections Thesis advisors: J.P. Murre, C.A.M. Peters (Institut Fourier, France)

September 24, 1997  Leiden University

A.K. Sinha Estimating failure probability when failure is rare: multidimensional case Thesis advisor: L. de Haan

October 2, 1997  Erasmus University
M. Maes  Mathematical Methods for Reflector Design

Thesis advisor: P.W. Hemker

October 6, 1997  University of Amsterdam

M. Fiocco Statistical Estimation for the Supercritical Contact Process Thesis advisor: W.R. van Zwet

October 15, 1997  Leiden University

M. Dekker Moduli spaces of stable sheaves on abelian surfaces Thesis advisor: G. van der Geer

November 17,1997  University of Amsterdam

J.J.B. de Swart Parallel software for implicit differential equations Thesis advisor: P.J. van der Houwen

November 28, 1997  University of Amsterdam
A.B. Ermakov  Percolation and Coalescing Particle Systems

Thesis advisor: M.S. Keane

December 1, 1997  Delft University of Technology

E. de Klerk Interior Point Methods for Semidefinite Programming Thesis advisor: F.A. Lootsma

December 8, 1997  Delft University of Technology

Number of Researchers

Situation at July 1, 1997

I. Funded by universities  II. Funded by NWO Total
& KNAW  I + II
perma- nent {post \atop doc} Ph.D. total {post \atop doc} Ph.D. total

UvA 19  1  11  31  4  12  16  47
VU  15  1  9  25  -  -  -  25
TUD 27  -  5  32  1  3  4  36
TUE 8  -  2  10  -  -  -  10
RUL 18  2  12  32  5  3  8  40
EUR 11  -  1  12  -  1  1  13
Total 98  4  40  142  10  19  29  171
CWI participants  18

31

Publications

(NOTE: Some publications who were missing in the Biennual Report 1995 & 1996 are also included)

Algebra and Geometry

1.1. Number Theory

Papers in journals and proceedings

H. Boender, H.J.J. te Riele, Factoring integers with large prime variations of the quadratic sieve, Experiment. Math 5 (1996), 257-273. W. Bosma, J. Cannon, C. Playoust, The Magma Algebra System I: The User Language, J. Symbolic Computation 24 (1997), 235-265. W. Bosma, J. Cannon, A. Steel, Lattices of compatibly embedded finite fields”, J. Symbolic Computation 24 (1997), 351-369. A. Bremner, R.J. Stroeker, N. Tzanakis, On sums of consecutive squares, J. Number Th. 62 (1997), 39-70. G.L. Cohen, H.J.J. te Riele, Iterating the Sum-of-Divisors Function, Experimental Math. 5(1996) 91–100. Erratum in 6 (1997), 177. K. Dajani, C. Kraaikamp, On approximation by Lüroth Series, J. Théor. Nombres Bordeaux 8 (1996), 331-346. J.-M. Deshouillers, G. Effinger, H. te Riele, D. Zinoviev, A Complete Vinogradov 3-Primes Theorem Under the Riemann Hypothesis, Electronic Research Announcements AMS, 3 (1997), 99–104. J.-H. Evertse, The number of solutions of the Thue-Mahler equation, J. reine angew. Math. 482 (1997), 121-149. J.-H. Evertse, The number of algebraic numbers of given degree approximating a given algebraic number, Analytic Number Theory, proc. conf. Kyoto 1996, Y. Motohashi, ed., pp. 53-84. London Math. Soc. Lecture Note Ser. 247, Cambridge University Press, 1997.

J.-H. Evertse, K. Győry, The number of families of solutions of decomposable form equations, Acta Arith. 80 (1997), 367-394. P. Montgomery, S. Cavallar, H. te Riele, A new world record for the special number sieve factoring method, CWI Quarterly 10 (1997), 105–107. T.N. Shorey and R. Tijdeman, Irrationality criteria for numbers of Mahler’s type, Analytic Number Theory, ed. by Y. Motohashi, LMS Lecture Note Series 247, Cambridge University Press, pp. 343-351. B. de Smit, A differential criterion for complete intersections, Collect. Math 48 (1997), 85–96. B. de Smit, The different and differentials for local fields with imperfect residue fields, Proc. Edinburgh Math. Soc. 40 (1997), 353–365. B. de Smit, H. W. Lenstra, Jr., Explicit construction of universal deformation rings, Modular Forms and Fermat’s Last Theorem (Cornell, Silverman, Stevens, Eds), pp. 313–326, Springer-Verlag 1997. B. de Smit, H. W. Lenstra, Jr., Finite complete intersection algebras and the completeness radical, J. Algebra 196 (1997), 520–531. B. de Smit, K. Rubin, R. Schoof, Criteria for complete intersections, Modular Forms and Fermat’s Last Theorem (Cornell, Silverman, Stevens, Eds), Springer-Verlag 1997. P. Stevenhagen, H. W. Lenstra, Jr., Class field theory and the first case of Fermat’s Last Theorem, Modular Forms and Fermat’s Last Theorem (Cornell, Silverman, Stevens, Eds), pp. 343–355, Springer-Verlag, 1997. P. Stevenhagen, P. Moree, Prime divisors of Lucas sequences, Acta Arith. 82 (1997), 403–410. C.L. Stewart and R. Tijdeman, On the greatest prime factor of (ab + 1)(ac + 1)(bc + 1), Acta Arith. 97 (1997), 93-101.

Ph.D. Theses

H. Boender, Factoring Large Integers with the Quadratic Sieve, Leiden University, 1997. M. Elkenbracht-Huizing, Factoring Integers with the Number Field Sieve, Leiden University, 1997.

Reports and pre-prints

G. Barat, R.F. Tichy, R. Tijdeman, Digital blocks in linear numeration systems, preprint.

F. Beukers, T.N. Shorey, R. Tijdeman, Irreducibility of polynomials and arithmetic progressions with equal products of terms, preprint. N. Bruin, The diophantine equations x^{2} ± y^{2} = ±z^{6} and x^{2} + y^{8} = z^{3}, Report W-97-21, Leiden University. J.-M. Deshouillers, H.J.J. te Riele, Y. Saouter, New experimental results concerning the Goldbach conjecture, preprint. J.-H. Evertse, The number of solutions of linear equations in roots of unity, preprint. J.-H. Evertse, H.P. Schlickewei, The Absolute Subspace Theorem and linear equations with unknowns from a multiplicative group, preprint. A. Gee, P. Stevenhagen, Generating class fields using Shimura reciprocity, WINS report 97-15, Univ. Amsterdam.

A. Heinis, R. Tijdeman, Extended balanced and stiff words, Report W-97-15, Leiden University. H. te Riele, Computational sieving applied to some classical numbertheoretic problems, preprint. B. de Smit, On Brauer relations for S-class numbers, WINS report 97-10, Univ. Amsterdam. R.J. Stroeker, N. Tzanakis, On the elliptic logarithm method for elliptic diophantine equations: Reflections and an improvement, Report 9739/B Econometric Institute, Erasmus University Rotterdam (1997), 26 pp. R.J. Stroeker, B.M.M. de Weger, Elliptic binomial diophantine equations, Report 9723/B Econometric Institute, Erasmus University Rotterdam (1997), 26 pp.

Other publications

W. Bosma, P. Stevenhagen, Priemgetallenserie, Pythagoras 1997-1998. P. Stevenhagen, Driehoeken en vierkanten, Pythagoras 5 (1997).

1.2. Algebraic and Analytic Geometry

Papers in journals and proceedings

E. Amerik, On codimension-two subvarieties of hypersurfaces, J. reine und angew. Mathematik 483 (1997), 61-73 J. Bochnak, W. Kucharz, On homology classes represented by real algebraic varieties, Singularities Symposium, Banach Center, Warsaw 1997 , p. 7-22. M. Boguslavsky, On the number of solutions of polynomial systems. Finite Fields and their Applications, 3 (1997), 287-299. R. Dijkgraaf, Chiral Deformations of Conformal Field Theories Nucl. Phys., B493 [FS] (1997) 588-612. R. Dijkgraaf, G. Moore, Balanced Topological Field Theories, Commun. Math. Phys., 185 (1997) 411-440. R. Dijkgraaf, G. Moore, E. Verlinde and H. Verlinde, Elliptic Genera of Symmetric Products and Second Quantized Strings, Commun. Math. Phys., 185 (1997) 197-209. R. Dijkgraaf, E. Verlinde and H. Verlinde, BPS Spectrum of the Five-Brane and Black Hole Entropy, Nucl. Phys., B486 (1997) 77-88. R. Dijkgraaf, E. Verlinde and H. Verlinde, BPS Quantization of the Five-brane, Nucl. Phys., (1997) 89-113. R. Dijkgraaf, E. Verlinde and H. Verlinde, Counting Dyons in N = 4 String Theory, Nucl. Phys., B484 (1997) 543. R. Dijkgraaf, E. Verlinde and H. Verlinde, Matrix String Theory, Nucl. Phys., B500 (1997) 43-61. R. Dijkgraaf, E. Verlinde and H. Verlinde, 5D Black Holes and Matrix Strings, Nucl.Phys., B506 (1997) 121-142. G. van der Geer, M. van der Vlugt, How to construct curves over finite fields with many points, Arithmetic Geometry, (Ed. F. Catanese), pp 169-189, Cambridge University Press 1997. J. van Hamel, Real algebraic cycles on complex projective varieties, Math. Zeitschrift, 225, pp 177-198 (1997) N.P. Landsman, Poisson spaces with a transition probability Rev. Math. Phys., 9 (1997), 29-57. N.P. Landsman, Classical reduction and quantum induction in constrained systems Proc. XXI Int. Colloquium on Group Theoretical Methods in Physics, Goslar 1996, Vol. 1, eds. H.-D. Doebner, W. Scherer, and P. Nattermann (World Scientific, Singapore), pp. 368-372 (1997). N.P. Landsman and K.K. Wren, Constrained quantization and θ-angles Nucl. Phys., B502 [PM] (1997), 537-560.

C.O.B. de Mirleau, Combinatorical aspects of the Schwinger-Dyson equation, Journal of Geometry and Physics, vol. 21, no. 4 (1997) 357-380. J.F. Morales and M. Serone, BPS states and supersymmetric index in N = 2 type I string vacua, Nucl. Phys., B501 (1997) 427. J. Nagel, The Abel-Jacobi map for complete intersections, Indag. Math 8 (1997), 95-113. C.D.D. Neumann, Perturbative BPS Algebras in Superstirng Theory, Nucl. Phys., B499 (1997) 596-620. Ch. Okonek, A. van de Ven, Cubic forms and complex 3-folds, l’Enseignement Mathématique, 41, 297-333 (1995) M. Serone, N = 2 Type I-Heterotic duality and higher derivative F-terms, Phys. Lett., B395 (1997) 42. M. van der Vlugt, Surfaces and the weight distribution of a family of codes. IEEE Trans. on Info. Theory 43 (1997), 1354-1360.

Ph.D. Theses

E. Amerik, On the geometry of smooth hypersurfaces in projective space Leiden University, 1997. M. Dekker, Moduli spaces of stable sheaves on abelian surfaces, University of Amsterdam, 1997. J. van Hamel, Algebraic cycles and topology of real algebraic varieties, Free University Amsterdam, 1997. C.O.B. de Mirleau, Foundations of functional integration and the Schwinger-Dyson equation, University of Amsterdam, 1997. J.F.A.A. Nagel, The image of the Abel-Jacobi map for complete intersections, Leiden University, 1997. P.G.J. Nijsse, Stable vector bundles on quintic surfaces, Leiden University, 1997. C.M. Schuhmann, Morphisms between Fano threefolds, Leiden University, 1997.

Reports and pre-prints

J. Bochnak, M. Coste-Roy, Real algebraic geometry, 450 pp. To appear in: Ergebnisse der Math. J. Bochnak, W. Kucharz, On dividing real algebraic varieties, To appear in Math. Proc. Cambr. Philo. Soc.

J. Bochnak, W. Kucharz, The Weierstrass Approximation theorem and a characterization of the unit circle, To appear in: Proc. AMS J. Bochnak, W. Kucharz, R. Silhol, Morphisms, line bundles and moduli spaces in real algebraic geometry, To appear in: Publ. Math. IHES 86 R. Dijkgraaf, Les Houches Lectures on Fields, Strings and Duality, hepth/9703136. R. Dijkgraaf, E. Verlinde and H. Verlinde, Notes on Matrix and Micro Strings, hep-th/970401. G. van der Geer, The Chow Ring of the Moduli Space of Abelian Threefolds, Preprint February 1997. alg-geom/9702009. To appear in Journal of Algebraic Geometry. G. van der Geer, M. van der Vlugt, Tables for the function N_{q}(g). G. van der Geer, M. van der Vlugt, Generalized Reed-Muller codes and curves with many points, Report W-97-22, Leiden University. G. van der Geer, M. van der Vlugt, Constructing curves over finite fields with many points by solving linear equations, Report W-97-29, Leiden University. E. Knoops, M. van der Vlugt, On generalized Hamming weights of BCH(2) dual, Report W97-14. Leiden University 1997. N.P. Landsman, Constrained quantization in algebraic field theory, to appear in Proc. XII Int. Congress of Mathematical Physics (Brisbane 1997), eds. A.J. Bracken and D. de Wit. N.P. Landsman, Mathematical Topics between Classical and Quantum Mechanics, 580p., Springer-Verlag New York, to appear (1998). N.P. Landsman, Simple new axioms for quantum mechanics, To appear in: Int. J. Theor. Phys. 37 (1998), 343-348. C.O.B. de Mirleau, Further illustration of the use of the Frobenius-Schwinger-Dyson equations, hep-th/9703061. J.F. Morales, C.A. Scrucca and M. Serone, A note on supersymmetric D-brane dynamics, hep-th/9709063 M. Serone, A comment on the R^{4}-coupling in (M)atrix theory, hep-th/9711031. R. Re, Multiplication of sections and Clifford bounds for stable vector bundles on curves, Report 1997; To appear in Communications in Algebra. R. Re, Some remarks on vector spaces of nilpotent matrices, Report 1997.

Other publications

R. Dijkgraaf et al., String Theory, Gauge Theory and Quantum Gravity, Proceedings Trieste Spring School and Workshop 1996, Eds. R. Dijkgraaf et al. Nucl. Phys. B (Proc. Suppl.) (1997). G. van der Geer, Knoten, Brocker Festschrift.

1.3. Topology and Dynamical Systems

Papers in journals and proceedings

J. Baars, H. Gladdines, On the linear invariance of Lindelöf numbers, Canad. Math. Bull. 39 (1996), 129-137 (appeared in 1997) A. Bella and J. van Mill, Tight points and countable fan-tightness, Top. Appl. 76 (1997) 1-7. W. W. Comfort and J. van Mill, How many ω-bounded subgroups?, Top. Appl. 77 (1997) 105-113. J.J. Dijkstra, A dimension raising hereditary shape equivalence, Fundamenta Mathematicae 149 (1996) 265-274 (appeared in 1997) J. J. Dijkstra and J. van Mill, Projections of planar Cantor sets in potential theory, Indag. Math. 8 (1997) 173–180. J. J. Dijkstra and J. van Mill, Two point set extensions – a counterexample, Proc. Amer. Math. Soc. 125 (1997), 2501-2502 J. J. Dijkstra and J. van Mill, On remainders without arcs, Bull. Polon. Acad. Sci. Sér. Math., Astronom. Phys., Phys. 44 (1996) 263– 266. (published in 1997) J. J. Dijkstra and J. van Mill, Groups without convergent sequences, Top. Appl. 74 (1996) 275–282. A. Dow, and K.P. Hart, A new subcontinuum of βR \ R, Proceedings of the American Mathematical Society (1997) 1861–1871 G. Levin and S.J. van Strien, Total disconnectedness of the Julia set of real polynomials, Asterisque (1997). W. Marciszewski, On topological embeddings of linear spaces, Mathematische Annalen 308 (1997) 21-30 W. Marciszewski, A function space C_{p}(X) not linearly homeomorphic to C_{p}(X) × R, Fundamenta Mathematicae 153 (1997) 125-140 W. Marciszewski, On hereditary Baire products, Bul. Polish Academy Sci. 45 (1997), 247-250 W. Marciszewski and J. Pelant, Absolute Borel sets and function spaces, Trans. Amer. Math. Soc. 349 (1997) 3585-3596

Reports and pre-prints

G. Levin and S.J. van Strien, Total disconnectedness of the Julia set of real polynomials, preprint. G. Levin and S.J. van Strien, Bounds for interval maps with one inflection point II, preprint (1997).

Other publications

T. Koetsier and J. van Mill, General topology, in particular dimension theory, in the Netherlands: the decisive influence of Brouwer’s intuitionism, Handbook of the history of general topology (C. E. Aull and R. Lowen, eds.), Kluwer Academic Publishers, Dordrecht, 1997, pp. 135–180.

1.4. Theoretical Foundations of Computer Sciences

Papers in journals and proceedings

M.H. ter Beek, Teams in Grammar Systems: Sub-context-free Cases, Lecture Notes in Computer Science, 1218 (1997) 197-216. R. Bloem, J. Engelfriet, Monadic Second Order Logic and Node Relations on Graphs and Trees, Lecture Notes in Computer Science, 1261 (1997) 144-161. H.L. Bodlaender, J. Engelfriet, Domino Treewidth, Journal of Algorithms, 24(1997), 94-123. J. Dassow, G. Paun, G. Rozenberg, Grammar systems, In: G. Rozenberg, A. Salomaa (eds.), Handbook of Formal Languages, Volume 2: Linear Modelling: Background and Application, pages 155-213, Springer-Verlag, 1997. A. Ehrenfeucht, T. Harju, G. Rozenberg, 2-structures - A Framework for Decomposition and Transformation of Graphs, In: G. Rozenberg (ed.), Handbook of Graph Grammars and Computing by Graph Transformation, Volume 1: Foundations, pages 401-478. World Scientific, 1997. A. Ehrenfeucht, T. Harju, G. Rozenberg, Invariants of Inversive 2-Structures on Groups of labels, Mathematical Structures in Computer Science, 7 (1997) 303-327.

A. Ehrenfeucht, G. Rozenberg, K. Salomaa, Semantics of Nonsequential Tree-Based Computation Schemes, Fundamenta Informaticae, 29 (1997) 305-324. J. Engelfriet, Context-Free Graph Grammars, In: G. Rozenberg, A. Salomaa (eds.), Handbook of Formal Languages, Volume 3: Beyond Words, pages 125-213, Springer-Verlag, 1997. J. Engelfriet, V. van Oostrom, Logical Description of Context-Free Graph Languages, Journal of Computer and System Sciences, 55 (1997) 489-503. J. Engelfriet, G. Rozenberg, Node Replacement Graph Grammars, In: G. Rozenberg (ed.), Handbook of Graph Grammars and Computing by Graph Transformation, Volume 1: Foundations, pages 1-94. World Scientific, 1997. J. Engelfriet, J.-J. Vereijken, Context-free Graph Grammars and Concatenation of Graphs, Acta Informatica, 34 (1997) 773-803. R. Freund, G. Paun, G. Rozenberg, A. Salomaa, Watson-Crick Finite Automata, In: 3rd DIMACS Workshop on DNA Based Computers, pages 305-317, 1997. H.J. Hoogeboom, A. Muscholl, The Code Problem for Traces - Improving the Boundaries (Note), Theoretical Computer Science, 172 (1997) 309-321. H.J. Hoogeboom, P. ten Pas, Monadic Second-Order Definable Text Languages, Theory of Computing Systems, 30 (1997) 335-354. L. Kari, G. Rozenberg, A. Salomaa, L systems, In: G. Rozenberg, A. Salomaa (eds.), Handbook of Formal Languages, Volume 1: Word, Language, Grammar, pages 253-328, Springer-Verlag, 1997. N.W. Keesmaat, H.C.M. Kleijn, Net-based Control versus Rational Control, The Relation between ITNC Vector Languages and Rational Languages, Acta Informatica, 34 (1997) 23-57. N.W. Keesmaat, H.C.M. Kleijn, Restrictions and Representations of Vector Controlled Concurrent Systems Behaviours, Theoretical Computer Science, 179 (1997) 61-102. A. Mateescu, G.R. Mateescu, G. Rozenberg, A. Salomaa, Shufflelike Operations on Omega-Words, Lecture Notes in Computer Science, 1218 (1997) 395-411. A. Mateescu, G. Rozenberg, A. Salomaa, Geometric Transformations of Language Families: The Power of Symmetry, International Journal of Foundations of Computer Science, 8 (1997) 1-14.

A. Mateescu, G. Rozenberg, A. Salomaa, Syntactic and Semantic Aspects of Parallellism, Lecture Notes in Computer Science, 1337 (1997) 79-105. G. Paun, G. Rozenberg, Contextual Grammars and Formal Languages, In: G. Rozenberg, A. Salomaa (eds.), Handbook of Formal Languages, Volume 2: Linear Modelling: Background and Application, pages 215-235, Springer-Verlag, 1997. G. Paun, G. Rozenberg, A. Salomaa, Computing by Splicing: Programmed and evolving splicing systems, In: 1997 IEEE International Conference on Evolutionary Computation (ICEC’97), pages 273-278, 1997.

Reports and pre-prints

R. Bloem, J. Engelfriet, Characterization of Properties and Relations defined in Monadic Second Order Logic on the Nodes of Trees, Department of Computer Science, Leiden University, Technical Report No. TR97-03, 1997. A. Ehrenfeucht, T. Harju, P. ten Pas, G. Rozenberg, Permutations, Parenthesis Words, and Schröder Numbers, Leiden University, Technical Report No. TR97-04, 1997. R. Freund, G. Paun, G. Rozenberg, A. Salomaa, Watson-Crick Automata, Leiden University, Technical Report No. TR97-13, 1997. A. Ehrenfeucht, J. Hage, T. Harju, G. Rozenberg, Complexity Problems in Switching Classes of Graphs, Leiden University, Technical Report No. TR97-15, 1997. J. Hage, T. Harju, The Size of Switching Classes with Skew Gains, Department of Computer Science, Leiden University, Technical Report No. TR97-02, 1997. H.J. Hoogeboom, N. van Vugt, The power of H systems: does representation matter?, Report No. TR97-16, 1997.

Other publications

C. Fernandez, A. Merceron, V. Parada, G. Rozenberg, M. Solar (eds.), Petri Nets, Sciences, 1997. J. Mycielski, G. Rozenberg, A. Salomaa (eds.), Structures in Logic and Computer Science, A Selection of Essays in Honor of A. Ehrenfeucht, Lecture Notes in Computer Science 1261, Springer-Verlag, 1997.

G. Rozenberg (ed.), Handbook of Graph Grammars and Computing by Graph Transformation, Volume 1: Foundations, World Scientific, 1997. G. Rozenberg, A. Salomaa (eds.), Handbook of Formal Languages, Springer-Verlag, 1997, 3 volumes. G. Rozenberg, A. Salomaa (eds.), Special issue on Formal Language Theory, Theoretical Computer Science, Volume 183, 1997.

Analysis

2.1. Operator Theory and Applications

Papers in journals and proceedings

D.Z. Arov, M.A. Kaashoek, D.R. Pik, Minimal and optimal linear discrete time-invariant dissipative scattering systems, Integral Equations and Operator Theory 29 (1997), 127-154. J.A. Ball, I. Gohberg and M.A. Kaashoek, Nudelman interpolation and the band method, Integral Equations and Operator Theory 27 (1997), 253-284. J.-B. Baillon, Ph. Clément, A. Greven and F. den Hollander, On the attracting orbit of a nonlinear transformation arising from renormalization of hierarchically interacting diffusions, Part II: the noncompact case, J. Funct. Anal. 146 (1997), 236-298. H. Bart, T. Ehrhardt and B.Silbermann, Logarithmic residues, generalized idempotents, and sums of idempotents in Banach algebras, Integral Equations and Operator Theory, 29 (1997), 155-186. H. Bart, I. Gohberg, and M.A. Kaashoek, Wiener-Hopf equations and linear systems, Proc. Symposia Appl. Math., Amer. Math. Soc. Providence RI, 1996, pp. 115-128. Y. Bolshakov, C.V.M. van der Mee, A.C.M. Ran, B. Reichstein, L. Rodman, Erratum to ”Polar decompositions in finite dimensional indefinite scalar product spaces: special cases and applications.” (OT 87, Birkhäuser, Basel, 1996, 61-94) Integral Equations and Operator Theory 27 (1997), 497-501. Y. Bolshakov, C.V.M. van der Mee, A.C.M. Ran, B. Reichstein, L. Rodman, Extensions of isometries in finite dimensional indefinite scalar product spaces, SIAM Journal Matrix Analysis and Applications 18 (1997), 752-774.

Y. Bolshakov, C.V.M. van der Mee, A.C.M. Ran, B. Reichstein, L. Rodman, Polar decompositions in finite dimensional indefinite scalar product spaces: general theory, Linear Algebra and its Applications 261 (1997), 91-141. P.G. Dodds, T.K. Dodds, B. de Pagter, F.A. Sukochev, Lipschitz continuity of the absolute value and Riesz projections in symmetric operator spaces, J. Funct. Anal. 148 (1997), 28-69. C. Foias, A. Frazho, I. Gohberg, and M.A. Kaashoek, A timevariant version of the commutant lifting theorem and nonstationary interpolation, Integral Equations and Operator Theory 28 (1997), 158-190. C. Foias, A. Frazho, I. Gohberg, and M.A. Kaashoek, Parametrization of all solutions of the three chains completion problem, Integral Equations and Operator Theory 29 (1997), 455-490. G. Fusco and S.M. Verduyn Lunel, Order structures and the heat equation, J. Differential Eqns. 139 (1997), 104-145. I. Gohberg, M.A. Kaashoek and F. van Schagen, Operator blocks and quadruples of subspaces: classification and the eigenvalue completion problem, Lin. Alg. Appl. 269 (1997), 65-89. J.J. Grobler and C.B. Huijsmans, Doubly Abel bounded operators with single spectrum, Quaestiones Math. 18 (1995), 397–406. C.B. Huijsmans, A characterization of complex lattice homomorphisms on Banach lattice algebras, First International Conference in Abstract Algebra (Kruger Park, 1993), Quaestiones Math. 18 (1995), 131–140. C.B. Huijsmans, Disjointness preserving operators on Banach lattices, Operator theory in function spaces and Banach lattices, 173–189, Oper. Theory Adv. Appl., 75, Birkhäuser, Basel, 1995. C.B. Huijsmans and W.A.J. Luxemburg, On some aspects of the theory of lattice-ordered algebras, Lattice theory and its applications (Darmstadt, 1991), 103–120, Res.Exp. Math., 23, Heldermann, Lemgo, 1995. C.B. Huijsmans and A.W. Wickstead, Bands in lattices of operators, Proc. Amer. Math. Soc. 124 (1996), 3835–3841. M.A. Kaashoek, C.V.M. van der Mee and A.C.M. Ran, Weighting operator patterns of Pritchard-Salamon realizations, Integral Equations and Operator Theory 27 (1997), 48-70 L. Kérchy en J.M.A.M. van Neerven, Polynomially bounded operators whose spectrum on the unit circle has measure zero, Acta Sci. Math. (Szeged) 63 (1997), 559-570.

L. Lerer, A.C.M. Ran, J-spectral and J-inner/outer factorizations for a class of matrix polynomials, Integral Equations and Operator Theory 29 (1997), 23-51. J.M.A.M. van Neerven, Elementary operator-theoretic proof of Wiener’s Tauberian theorem, Rendic. Istit. Matem. Univ. Trieste, Suppl. Vol. XXVIII (1997), 279-284. B. de Pagter, W.J. Ricker, Boolean algebras of projections and resolutions of the identity of scalar-type spectral operators, Proc. Edinburgh Math. Soc. (2) 40 (1997), 425-435. B. de Pagter, A.R. Schep, Diagonals of positive semigroups, Integral Equations Operator Theory 27 (1997), 446-472. A.C.M. Ran and L. Rodman, Rate of stability of solutions of matrix polynomial and quadratic equations, Integral Equations and Operator Theory 27 (1997), 71-102. A.C.M. Ran and P. Zizler, On selfadjoint matrix polynomials with constant signature. Linear Algebra and its Applications 259 (1997), 133-153. Ph. Thijsse, Spectral Criteria for Complementary Triangular Forms, Integral Equations Operator Theory 27 (1997), 228-251. Ph. Thijsse, Upper Triangular Similarity of Upper Triangular Matrices, Linear Algebra Appl. 260, 119-149 (1997). S.M. Verduyn Lunel and D.V. Yakubovich, A functional model approach to linear neutral functional differential equations, Integral Equations and Operator Theory 27 (1997), 347-378.

Reports and pre-prints

Ph. Clément, G. Gripenberg and S.-O. Londen, Schauder estimates for equations with fractional derivatives, Helsinki University of Technology, Institute of Mathematics Research Reports A377, 27 pp., February 1997. Ph. Clément and S. Guerre-Delabrière, On the regularity of abstract Cauchy problems of order one and boundary value problems of order two, Publications Mathématiques de l’Université Pierre et Marie Curie Nr. 119, 44 pp., October 1997. A. Fischer en J.M.A.M. van Neerven, Robust stability of C_{0} -semigroup and an application to stability of delay equations, Institut für Dynamische Systeme, Report Nr. 413, 16 pp., September 1997.

A.C.M. Ran and L. Rodman, Stability of invariant subspaces of matrices with applications, Rapport no WS-488, Faculteit Wiskunde en Informatica, Vrije Universiteit.

2.2. Lie Groups, Special Functions, Approximation Theory and Complex Analysis

Papers in journals and proceedings

H. Bavinck, A difference operator of infinite order with Sobolev-type Charlier polynomials as eigenfunctions, Indag. Math. N.S. 7 (1996), 281-291. H. Bavinck, Linear perturbations of differential or difference operators with polynomials as eigenfunctions, J. Comput. Appl. Math. 78 (1997), 179-195; 83 (1997), 145-146. Yu. Bespalov and B. Drabant, Hopf (bi-)modules and crossed modules in braided monoidal categories, J. Pure Appl. Algebra 123 (1998), 105-129. M.G. de Bruin, K.G. Ivanov and A. Sharma, A conjecture of Schoenberg, Communications in the Analytic Theory of Continued Fractions 6 (1997), 103-105. M.G. de Bruin and C. Kamminga, Normalization procedures and Shannon’s entropy measure, Proc. Eighteenth Symposium on Information Theory in the Benelux, Veldhoven 1997, pp. 131-141. M.G. de Bruin and A. Sharma, Overconvergence of some simultaneous Hermite-Padé interpolants, Ann. Numer. Math. 4 (1997), 239-259. A. Bultheel, P. González-Vera, E. Hendriksen and O. Njåstad, A Favard theorem for orthogonal rational functions with poles on the unit circle, East J. Approx. 3 (1997), 21-37. A. Bultheel, P. González-Vera, E. Hendriksen and O .Njåstad, Orthogonal rational functions and nested disks, J. Approx. Theory 89 (1997), 344-371. A. Bultheel, P. González-Vera, E. Hendriksen and O. Njåstad, Rates of convergence of multipoint rational approximants and quadrature formulas on the unit circle, J. Comput. Appl. Math. 77 (1997), 77-102. A. Bultheel, P. González-Vera, E. Hendriksen and O. Njåstad, A rational moment problem on the unit circle, Methods Appl. Anal., 4(3) (1997), 283-310. G. van Dijk, Canonical representations, Vestnik Tambov Univ. 2 (1997), no.4, 350-366.

G. van Dijk and S.C. Hille, Canonical representations related to hyperbolic spaces, J. Funct. Anal. 147 (1997), 107-139. G. van Dijk and S.C. Hille, Maximal degenerate representations, Berezin kernels and canonical representations, in: Lie Groups and Lie Algebras, Their Representations, Generalizations, and Applications (eds. B. Komrakov, J. Krasil’shchik, G. Litvinov and A. Sossinsky), Kluwer Academic, Dordrecht, 1997. G. van Dijk and K. Mokni, Harmonic analysis on Gel’fand pairs associated with hyperbolic spaces, Russian J. Math. Phys. 5 (1997), no.2, 1-13. H.T. Koelink and P.G.A. Floris, A commuting q-analogue of the addition formula for disk polynomials, Constr. Approx. 13 (1997), 511-535. H.T. Koelink, Addition formulas for q-special functions, in: Special Functions, q-Series and Related Topics (eds. M.E.H. Ismail, D.R. Masson, M. Rahman), Fields Institute Communications 14, Amer. Math. Soc., 1997, pp. 109-129. H.T. Koelink, Convolution and addition theorems for q-orthogonal polynomials, in: Proceedings of the International Workshop on Orthogonal Polynomials in Mathematical Physics 1996 (eds. M. Alfaro, R. Álvarez-Nodarse, G. López Lagomasino and F. Marcellán), Madrid, 1997, pp. 113-122. J. Koekoek and R. Koekoek, Finding differential equations for symmetric generalized ultraspherical polynomials by using inversion methods, in: Proceedings of the International Workshop on Orthogonal Polynomials in Mathematical Physics 1996 (eds. M. Alfaro, R. Álvarez-Nodarse, G. López Lagomasino and F. Marcellán), Madrid, 1997, pp. 103-111. T.H. Koornwinder, Special functions and q-commuting variables, in: Special Functions, q-Series and Related Topics (eds. M.E.H. Ismail, D.R. Masson and M. Rahman), Fields Institute Communications 14, Amer. Math. Soc., 1997, pp. 131-166. T.H. Koornwinder and N.M. Muller, The quantum double of a (locally) compact group, J. Lie Theory 7 (1997), 101-120. T.H. Koornwinder and A.M. Schwartz, Product formulas and associated hypergroups for orthogonal polynomials on the simplex and on a parabolic biangle, Constr. Approx. 11 (1997), 537-567. N.P. Landsman, Poisson spaces with a transition probability, Rev. Math. Phys. 9 (1997), 29-57. N.P. Landsman, Simple new axioms for quantum mechanics, Int. J. Theor. Phys. 37 (1998), 343-348.

N.P. Landsman, Classical reduction and quantum induction in constrained systems, in: Group 21, Vol. 1, (eds. H.-D. Doebner, P. Nattermann and W. Scherer), World Scientific, Singapore, 1997, pp. 368-372. N.P. Landsman and K.K. Wren, Constrained quantization and θ-angles Nucl. Phys. B 502 (1997), 537-560. F. Marcellán, H.G. Meijer, T.E. Pérez and M.A. Piñar, An asymptotic result for Laguerre-Sobolev orthogonal polynomials, J. Comput. Appl. Math. 87 (1997), 87-94. H.G. Meijer, Determination of all coherent pairs, J. Approx. Theory 89 (1997), 321-343. M.A. Monterie, Capacities of certain Cantor sets, Indag. Math., N.S. 8 (1997), 247-266. P.J. Oonincx, Analysing earthquake data: Tapered wavelet analysis, in: Proc. 2nd IEEE Symposium on Applications of Time-Frequency and Time-Scale Methods, 1997, pp. 13-16. E.M. Opdam and G.J. Heckman, Yang’s system of particles and Hecke algebras, Ann. of Math. (2) 145 (1997), 139-173; 146 (1997), 749-750. E.B. Saff and A.B.J. Kuijlaars, Distributing many points on a sphere, Math. Intelligencer 19 (1997), 5-11. J.V. Stokman, Multivariable big and little q-Jacobi polynomials, SIAM J. Math. Anal. 28 (1997), 452-480. J.V. Stokman, Two limit transitions involving multivariable Askey-Wilson polynomials, in: Quantum Groups and Quantum Spaces (eds. R. Budzynski, W. Pusz and S. Zakrzewski), Banach Center Publications 40 (1997), pp. 415-428. J.V. Stokman and T.H. Koornwinder, Limit transitions for BC type multivariable orthogonal polynomials, Can. J. Math. 49 (2) (1997), 373-404. N.M. Temme, Asymptotics and zero distribution of zeros of polynomials that are related to Daubechies wavelets, Appl. Comput. Harmon. Anal. 4 (1997), 414-428. N.M. Temme, Numerical algorithms for uniform Airy-type asymptotic expansions, Numer. Algorithms 15 (1997), 207-225. R.A. Zuidwijk, Quasicomplete factorizations of rational matrix functions, Integral Equations Oper. Theory 27 (1997), 111-124. R.A. Zuidwijk, Complementary triangular forms, Linear Operators, Banach Center Publications, Vol. 38, 1997, pp. 443-452.

R.A. Zuidwijk, Directional time-scale analysis with wavelet X-ray transform, in: Proceedings UK Symposium on Applications of Time-Frequency and Time-Scale Methods, University of Warwick, Coventry, 1997, pp. 169-171. R.A. Zuidwijk, The discrete and continuous wavelet X-ray transform, in: SPIE proceedings 3169, Wavelet Applications in Signal and Image Processing V, 1997, pp. 357-366.

Reports and pre-prints

H. Bavinck, Generalizations of Meixner polynomials which are eigenfunctions of a difference operator, Report 97-06, TUDelft Fac. Techn. Math. & Inf., 1997 to appear in: J. Difference Equ. Appl. H. Bavinck, Differential operators having Laguerre type and Sobolev type Laguerre polynomials as eigenfunctions: a survey, Report 97-07, TUDelft Fac. Techn. Math. & Inf., 1997 to appear in: “Special Functions and Differential Equations”, Proceedings Workshop Madras 1997. H. Bavinck, Differential and difference operators having orthogonal polynomials with two linear perturbations as eigenfunctions, Report 97-32, TUDelft Fac. Techn. Math. & Inf., 1997 to appear in: J. Comput. Appl. Math. P.R. Beneker, Strongly exposed points, Helson-Szegö weights and Toeplitz operators, Report 97-05, Math. preprint series, Univ. of Amsterdam, 1997 to appear in: Integral Equations Oper. Theory. K.A. Driver and N.M. Temme, Zero and pole distribution of diagonal Padé approximants to the exponential function, CWI Report MAS-R9701, 1997 to appear in: Quaest. Math. K.A. Driver and N.M. Temme, On polynomials related with Hermite-Padé approximations to the exponential function, CWI Report MAS-R9716, 1997 to appear in: J. Approx. Theory. K.A. Driver and N.M. Temme, Asymptotics and zero distribution of Padé polynomials associated with the exponential function, CWI Report MAS-R9726, 1997 to appear in: J. Comput. Appl. Math. G. van Dijk, Berezin kernels on tube domains, Preprint W 97-07 RUL, 1997. G. van Dijk and V.F. Molchanov, The Berezin form for rank one para-Hermitian symmetric spaces, Report W 97-05 RUL, 1997. G. van Dijk and V.F. Molchanov, Tensor products of maximal degenerate representations of the group SL(n, R), Report W 97-04 RUL, 1997.

G. van Dijk and A.A. Pasquale, Harmonic analysis on vector bundles over Sp(1, n)/Sp(1)timesSp(n), Report W 97-17 RUL, 1997. S.J.L. van Eijndhoven and P.J. Oonincx, Frames, Riesz systems and MRA in Hilbert spaces, CWI Report PNA-9701, 1997. H.T. Koelink and J. Van der Jeugt, Bilinear generating functions for orthogonal polynomials, Report 97-03 Math. preprint series, Univ. of Amsterdam, 1997. H.T. Koelink, Some basic Lommel polynomials, Report 97-07, Math. preprint series, Univ. of Amsterdam, 1997, to appear in: J. Approx. Theory. T.H. Koornwinder, F.A. Bais and N.M. Muller, Tensor product representations of the quantum double of a compact group, Report 97-14, Math. preprint series, Univ. of Amsterdam, 1997. F. Marcellán, H.G. Meijer, T.E. Pérez and M.A. Piñar, An asymptotic result for Laguerre-Sobolev orthogonal polynomials, Report 97-03, TUDelft Fac. Techn. Math. & Inf., 1997. H.G. Meijer, T.E. Pérez and M.A. Piñar, Asymptotics of Sobolev orthogonal polynomials of Laguerre-type, Report 97-45, TUDelft Fac. Techn. Math. & Inf., 1997. P.J. Oonincx, On time-frequency analysis and time-limitedness, CWI Report PNA-9720, 1997. Y.A. Sharshov, Spherical distributions on U(p, q) with respect to the subgroup U(1) × U(p − 1, q), Preprint W 97-08 RUL, 1997. Y.A. Sharshov and V.F. Molchanov, Representations of the universal covering group of the group SU(p, q) associated with a cone, Preprint W 97-09 RU, 1997. J.V. Stokman, On BC type basic hypergeometric orthogonal polynomials, Report 97-08, Math. preprint series, Univ. of Amsterdam, 1997. J.V. Stokman and T.H. Koornwinder, On some limit cases of Askey-Wilson polynomials, Report 97-14, Math. preprint series, Univ. of Amsterdam, 1997, to appear in: J. Approx. Theory. N.M. Temme, Analytical methods for a selection of elliptic singular perturbation problems, CWI Report MAS-R9727, 1997, to appear in: ”Recent Advances in Differential Equations”, proceedings of the First Pan-China Conference on Differential Equations. R.A. Zuidwijk, The wavelet X-ray transform, CWI Report PNA-9703, 1997.

Other publications

G. van Dijk, Thomas Stieltjes, Pythagoras 12 (1997), p.4. H.T. Koelink and T.H. Koornwinder, Addition theorems in the theory of special functions, Encyclopaedia of Mathematics Supplement Volume I (ed. M. Hazewinkel), Kluwer Acad. Publ., 1997, p.29. J.J.O.O. Wiegerinck, Bloch function, Encyclopaedia of Mathematics Supplement Volume I (ed. M. Hazewinkel), Kluwer Acad. Publ., 1997, p.130. J.J.O.O. Wiegerinck, BMO-space, Encyclopaedia of Mathematics Supplement Volume I (ed. M. Hazewinkel), Kluwer Acad. Publ., 1997, pp. 133-134. J.J.O.O. Wiegerinck, VMO-space, Encyclopaedia of Mathematics Supplement Volume I (ed. M. Hazewinkel), Kluwer Acad. Publ., 1997, p.480.

2.3. Non-lineair Partial Differential Equations

Papers in journals and proceedings

Ph. Clément, P. Felmer, E. Mitidieri, Homoclinic orbits for a class of infinite dimensional Hamiltonian systems. Annali della Scuola Normale Superiore di Pisa, Scienze Fisiche e Matematiche, 24, 4, 1997, p. 367-393. Ph. Clément, D.G. de Figueiredo, E. Mitidieri, Quasilinear elliptic equations with critical exponents. Topological Methods in Nonlinear Analysis, ISSN 1230-3429 (1996), 7, 1997, p. 133-170. Ph. Clément and G. Da Prato, Some results on stochastic convolutions arising in Volterra equations perturbed by noise, Rend. Atti Acc. Lincei 9 (1996), 147-153. Ph. Clément, R.C.A.M. van der Vorst, On the non-existence of homoclinic orbits for a class of infinite dimensional Hamiltonian systems. Proceedings of the American Mathematical Society, ISSN 0002-9939, 125, 4, 1997, p. 1167-1176. C.J. van Duijn, P. Knabner, Travelling wave behaviour of crystal dissolutionin porous media flow. Euro. Applied Mathematics 8, 49-72 (1997). C.J. van Duijn, R.E. Grundy, C.N. Dawson, Large time profiles in reactive solute transport. Transport in Porous Media 27, 57-84 (1997). V. A. Galaktionov, J. Hulshof & J.L. Vazquez, Extinction and focusing behaviour of spherical and annular flames described by a free boundary problem, J. Math. Pures et Appl. 76 (1997) 563-608.

V.A. Galaktionov & L.A. Peletier, Asymptotic behaviour near finite time extinction for the fast diffusion equation, Arch. Rational Mech. Anal. 139 (1997) 83-98. H.-Ch. Grunau, G. Sweers, Positivity for equations involving polyharmonic operators with Dirichlet boundary conditions. Mathematische Annalen, Springer-Verlag, 307, 1997, p. 589-626. H.-Ch. Grunau, G. Sweers, Classical solutions for some higher order semilinear elliptic equations under weak growth conditions. Nonlinear Analysis, Theory, Methods & Applications. Pergamon, ISSN 0362-546X (1996), 28, 5, 1997, p. 799-807. H.-Ch. Grunau, G. Sweers, The Maximum Principle and Positive Principal Eigenfunctions for Polyharmonic Equations. In: G. Caristi, E. Mitidieri; Proceedings of Reaction Diffusion Systems, Trieste, October 2-7, 1995, Marcel Dekker, Inc; ISBN 0-8247-0125-9, New York, USA, 1997, p. 163-182. H.-Ch. Grunau, G. Sweers, Positivity properties of elliptic Dirichlet problems of higher order, Nonlinear Analysis, T.M.A. 30 (1997), 5251-5258 (Proc. 2nd World Congress of Nonlinear Analysts). J. Hulshof, Selfsimilar solutions of Barenblatt’s model for turbulence, SIAM J. Math. Anal. 28 (1997) 33-48. J. Hulshof, Recent results on selfsimilar solutions of degenerate nonlinear diffusion equations, In: Reaction Diffusion Systems (eds. G. Caristi & E. Mitidieri), Lectures Notes in Pure and Applied Mathematics 194, Marcel Dekker, 1997. L.A. Peletier, Self-similar solutions of the second kind, Proceedings of the Ferrara meeting in honour of J. Serrin’s 65th birthday, (Eds. Pucci et al), Springer, Berlin, 1997. L.A. Peletier & W.C. Troy, Spatial patterns described by the Extended Fisher-Kolmogorov (EFK) equation: Periodic solutions, SIAM J. Math Anal. 28 (1997) 1317-1353. I. Peral and R.C.A.M. van der Vorst, On some quasilinear systems, Rocky Mount. Journ. 27(3) (1997) 913-927 FQ̇uirós, Asymptotic Behaviour of Nonlinear Diffusion Problems in Exterior Domains, FBP News 14 (1997) 6-7. G. Sweers, L^{n} is sharp for the anti-maximum principle. Journal of Differential Equations Academic Press, ISSN 0022-0396, 134, 1, 1997, p. 148-153.

Reports and pre-prints

H.W. Alt, C.J. van Duijn, A free boundary value problem involving a cusp. Part II: local analysis. CWI Report MAS-R9703. S.B. Angenent & R.C.A.M. van der Vorst, A superquadratic elliptic system and its Morse-Conley-Floer homology, Mathematical Institute Report 97-26, November 1997, to appear in Math. Zeitschrift. S.B. Angenent & R.C.A.M. van der Vorst, A priori bounds and renormalized Morse indices of solutions of an elliptic system, Mathematical Institute Report 97-27, November 1997. G.J.B. van den Berg, Uniqueness of solutions for the Extended Fisher-Kolmogorov equation, Mathematical Institute Report 97-24, November 1997. G.J.B. van den Berg & R.C.A.M. van der Vorst, A domain-wall between single-mode and bimodal states, Mathematical Institute Report 97-25, November 1997. Birindelli, E. Mitidieri and G. Sweers, Existence of the principal eigenfunction for cooperative elliptic systems in a general domain, preprint. C.M. Brauner, J. Hulshof & C. Schmidt-Laine, The saddle point property for focusing selfsimilar solutions in a free boundary problem, Mathematical Institute Report W97-12 (1997), to appear in: Proc. A.M.S. E.N. Dancer, D. Hilhorst, M. Mimura & L.A. Peletier, Spatial segregation limit of a competition-diffusion system, Leiden University Report W97-17. C.J. van Duijn, R.J. Schotting, Brine transport in porous media: On the use of Von Mises and similarity transformations. CWI Report MAS-R9724. C.J. van Duijn, G. Galiano, J.H. Knight, M.A. Peletier, How mangroves salinize the soil. CWI Report MAS-R9728. B.W. van de Fliert, The behaviour of the boundary in two Stefan problems with negative latent heat, preprint. B.W. van de Fliert & R. van der Hout, Stress-driven diffusion in a drying liquid paint layer, to appear in: E.J.A.M. B.W. van de Fliert & R. van der Hout, A nonlinear Stefan problem with negative latent heat, arising in a diffusion model, preprint. B. Franchi & L.A. Peletier, Ground states for Gaussian curvature type equations, Leiden University Report W97-11, to appear in Asymptotic Analysis.

P. Freitas, G. Sweers, Positivity results for a nonlocal elliptic equation, to appear in The Royal Society of Edinburgh, Proceedings A. V.A. Galaktionov, L.A. Peletier & J.-L. Vazquez, Asymptotics of the fast-diffusion equation with critical exponent, Mathematical Institute Report W97-18. H.-Ch. Grunau and G. Sweers, The role of positive boundary data in the generalized clamped plate equation, to appear in ZAMP. J. Hulshof & J.R. King, Asymptotic analysis of an elliptic-parabolic moving boundary problem, Mathematical Institute Report W97-23 (1997), to appear in: Appl. Math. Lett. J.Hulshof & A.E. Shishkov, The finite speed of propagation property for the thin film equation in terms of the L^{1}-norm, Mathematical Institute Report W97-02(1997), to appear in: Adv. Diff. Equ. J.F. Kaashoek, Pattern formation by potentialised partial differential equations, Econometric Institute, Rotterdam, Report 9751/B (20 pag). R. Manásevich, G. Sweers, A noncooperative system with p-Laplacians that reserves positivity, to appear in Nonlinear Analysis, T.M.A. V.J. Mizel, L.A. Peletier & W.C. Troy, Periodic phases in second order materials, Leiden University Report W97-13. V.J. Mizel, L.A. Peletier & W.C. Troy, Periodic phases in second order materials, and higher order model equations, to appear in the Proceedings of the meeting in Pont-à-Mousson, June 1997. L.A. Peletier, A.I. Rotariu-Bruma & W.C. Troy, Pulse-like spatial patterns described by higher order model equations, Leiden University Report W97-16, to appear in J. Diff. Equ.. L.A. Peletier, A.I. Rotariu-Bruma & W.C. Troy, A generalisation of the Extended Fisher-Kolmogorov equation with an application to pulse propagation along optical fibers, to appear in the Proceedings of the meeting in Pont-à-Mousson, June 1997. F. Quirós & J.L. Vázquez, Asymptotic convergence of the Stefan Problem to Hele-Shaw, to appear in: Trans. Amer. Math. Soc. C.J. Reinecke and G. Sweers, A positive solution on to a system of elliptic equations of Fitzhugh-Nagumo type, preprint. N. Stavrakakis, G. Sweers, Positivity for a noncooperative system of elliptic equations in R^{n}, to appear in Differential and Integral Equations.

R.J. Schotting, H. Moser, S.M. Hassanizadeh, High-concentration gradient dispersion: Experiments, analysis and approximation. CWI Report MAS-R9734. R.J. Schotting, J.F. Scheid, The interface between fresh and salt groundwater in heterogeneous aquifers: a numerical approach. CWI Report MAS-9735.

2.4. Numerical Analysis

Papers in journals and proceedings

E.F.F. Botta, K. Dekker, Y. Notay, A. van der Ploeg, C. Vuik, F.W. Wubs, P.M. de Zeeuw, How fast the Laplace equation was solved in 1995, J. Applied Numerical Mathematics (1997) 24, 439–455. J.C. Butcher and S. Tracogna, Order conditions for Two-Step Runge-Kutta methods, Appl. Numer. Math. 24, 351–364. O.F. Cools, G.C. Herman, R.M. Van der Weiden and F.B. Kets, Fast computation of the 3-D Radon transform, Geophysics (1997) 62, 362– 364. E.F. Ernst and G.C. Herman, Computation of Green’s function of laterally varying media by means of a modal expansion, in: Chavent, G. and Sabatier, P.C. (eds.), Inverse problems of wave propagation and diffraction, Berlin, Springer, ISBN 3-540-62865-7 (Lecture Notes in Physics; 486), 294– 305 (1997) . E.F. Ernst and G.C. Herman, Estimation of a dispersive background medium with application to field data, SEG 67th Annual Meeting, Expanded Abstract, L.J. Edleson, M. McCormack (eds.), Dallas, SEG, ISSN 1052-3812, (1997) 1460–1463. E.F. Ernst and G.C. Herman, Reconstruction of the near-surface phase velocity, 59th EAGE Conference, Genève, Extended Abstract, Feenstra D.J. (ed.), Geneva,1997, EAGE, ISBN 90-73834-04-X, pp. P029. J. Frank and W. Hundsdorfer, J.G. Verwer, On the stability of implicit-explicit linear multistep methods, Applied Numerical Mathematics 25, (1997) 193–205. P.W. Hemker, Finite volume multigrid for 3D-problems, Notes on Numerical Fluid Mechanics, 57, (1997) Vieweg, Braunschweig, 485-509. P.W. Hemker, B. Koren, W.M. Lioen, M. Nool and H.T.M. van der Maarel, Multigrid for steady gas dynamics problems, Notes on Numerical Fluid Mechanics 57, (1997) Vieweg, Braunschweig, 393–417.

P.W. Hemker, G.I.Shishkin, L.P.Shishkina, The use of defect correction for the solution of parabolic singular perturbation problems, ZAMM 77, (1997) 1, 59-74. P.W. Hemker and P.M. de Zeeuw, A data structure for 3-dimensional sparse grids, Notes on Numerical Fluid Mechanics 57, (1997) Vieweg, Braunschweig, 443–484. G.C. Herman, Waves generated by highs-speed trains, Congress SEG/DALLA 1997. Technical Program. Editors: McCormack, M., Edleson, L. Publisher: Society of Exploration Geophysicists, 1913–1916. (1997) ISSN 1052-3812. W. Hoffmann & J.J.B. de Swart, Approximating Runge-Kutta matrices by triangular matrices, BIT 37, (1997) 346-354. K.J. in ’t Hout, Stability analysis of Runge-Kutta methods for systems of delay differential equations, IMA J. Numer. Anal. 17, (1997) 17–27. P.J. van der Houwen, Parallel iterative linear solvers for Runge-Kutta discretizations of VIDEs, Proceedings of the 15th IMACS World Congress, ed. Achim Sydow, Wissenschaft und Technik Verlag, Berlin, Volume 2, (1997) P.J. van der Houwen and E. Messina, Parallel linear system solvers for Runge-Kutta-Nyström methods, J. Comput. Appl. Math. 82, (1997) 407-422. P.J. van der Houwen, B.P. Sommeijer, Splitting methods for threedimensional transport models with interaction terms, J. Scientific Computing 12, (1997) P.J. van der Houwen, B.P. Sommeijer, Euler-Chebyshev methods for integro-differential equations, Appl. Numer. Math. 24, (1997) 203-218. P.J. van der Houwen, B.P. Sommeijer, J. Kok, The iterative solution of fully implicit discretizations of three-dimensional transport models, Appl. Numer. Math. 25, (1997) 243–256. P.J. van der Houwen and J.J.B. de Swart, Parallel linear system solvers for Runge-Kutta methods, Adv. Comp. Math. 7, (1997) 157-181. P.J. van der Houwen and J.J.B. de Swart, Parallel linear system solvers for Runge-Kutta methods, Proceedings of the 15th IMACS World Congress, ed. Achim Sydow, Wissenschaft und Technik Verlag, Berlin, Volume 2, (1997) 63-68. P.J. van der Houwen and W.A. van der Veen, Waveform relaxation methods for implicit differential equations, Adv. Comp. Math. 7, (1997) 183-197.

B. Koren, P.W. Hemker and C.T.H. Everaars, Multiple semi-coarsened multigrid for 3D CFD, Proceedings 13th AIAA Computational Fluid Dynamics Conference, Snowmass Village, CO, (AIAA-paper 97-2029), (1997) American Institute of Aeronautics and Astronautics, Reston, VA, 892–902. B. Koren, P.W. Hemker and P.M. de Zeeuw, Semi-coarsening in three directions for Euler-flow computations in three dimensions, Notes on Numerical Fluid Mechanics 57, (1997) Vieweg, Braunschweig, 547–567. B. Koren, B. van Leer, Improving Euler computations at low Mach numbers, Notes on Numerical Fluid Mechanics 57, 419-441, (1997) Vieweg, Braunschweig. R. Lauterbach and J.A. Sanders, Bifurcation analysis for spherically symmetric systems using invariant theory J. of Dynamics and Differential Equations 9, (1997) 535–560. M. van Loon, A.W. Heemink, P.J.H. Builtjes, Data assimilation for atmospheric chemistry models, In: Measurements and Modelling in Environmental Pollution, eds. R. San Jose and C.A. Brebbia, Computational Mechanics Publications, Southhampton, UK, (1997) pp. 293–300. M.R.F. Massier, G.C. Herman and D.J. Verschuur, Removal of overburden transmission effects by deterministic deconvolution, Geophysics 62, (1997) 1612–1616. E. Messina, J.J.B. de Swart and W.A. van der Veen, Parallel iterative linear solvers for Multistep Runge-Kutta methods, J. Comput. Appl. Math. 85, (1997) 145-167. A.J.H. Muijres and G.C. Herman, The influence of spatial correlations on effective media, EAGE Congres, Geneva. Extended abstracts of papers, 2 pages. Editor: Feenstra D.J., EAGE, ISBN 90/73834-04-X. (1997) J.A. Sanders and J.P. Wang, Hodge decomposition and conservation laws, Mathematics and Computers in Simulation 44, (1997) 483–493. J.A. Sanders and J.P. Wang, Classification of conservations laws for KdV -like equations, Mathematics and Computers in Simulation 44, (1997) 471–481. A. Sandu, J.G. Verwer, J.G. Blom, E.J. Spee, G.R. Carmichael, Benchmarking stiff ODE solvers for atmospheric chemistry problems II: Rosenbrock solvers, Atmospheric Environment 31, (1997) 3459–3472. A. Sandu, J.G. Verwer, M. van Loon, G.R. Carmichael, F.A. Potra, D. Dabdub, J.H. Seinfeld, Benchmarking stiff ODE solvers for atmospheric chemistry problems I: Implicit versus explicit, Atmospheric Environment 31, (1997) 3151–3166. A.G.J. Sevink and G.C. Herman, Three-dimensional inverse scattering versus imaging methods, Journal of Seismic Exploration 6, (1997) 227-237. B.P. Sommeijer and J. Kok, Domain decomposition for an implicit shallowwater transport solver, in: B. Hertzberger, P. Sloot (eds.), Proc. HPCN Europe 1997 Conference, Lect. Notes in Comp. Sc. 1225, Springer, pp. 379–388. E.J. Spee, Development of advection schemes for global 3D chemistrytransport models, in: H. Power and T. Tirabassi and C.A. Brebbia (eds.), Proc. Air Pollution V: Modelling, Monitoring and Management, Southampton, (1997) pp. 177–186. M.N. Spijker, Numerical stability, resolvent conditions and delay differential equations, Appl. Numer. Math. 24, (1997) 233–246. M.N. Spijker and F.A.J. Straetemans, Error growth analysis via stability regions for discretizations of initial value problems, BIT 37, (1997) 442–464. J.J.B. de Swart, A simple ODE solver based on 2-stage Radau IIA, J. Comput. Appl. Math. 84, (1997) 277-280. J.J.B. de Swart and G. Söderlind, On the construction of error estimates for implicit Runge-Kutta methods, J. Comput. Appl. Math. 86, (1997) 347-358. J.G. Verwer and B.P. Sommeijer, Stability analysis of an odd-even hopscotch method for three-dimensional advection-diffusion problems, SIAM J. Numer. Anal. 34, (1997) 376–388. P.M. de Zeeuw, Multipe semi-coarsening techniques. Notes on Numerical Fluid Mechanics 57, Braunschweig, (1997) 511–545.

Ph.D. Theses

M. Maes, Mathematical Methods for reflector Design, University of Amsterdam, 1997. F.A.J. Straetemans, Stability and Resolvent Conditions for Discretizations of Initial Value Problems, Leiden University 1997. J.J.B. de Swart, Parallel software for implicit differential equations, University of Amsterdam, 1997. W.A. van der Veen, Parallelism in the numerical solution of ordinary implicit differential equations, University of Amsterdam, 1997.

P.M. de Zeeuw, Acceleration of Iterative Methods by Coarse Grid Corrections, University of Amsterdam, 1997.

Reports and pre-prints

F. Beukers, J.A. Sanders and J.P. Wang, One Symmetry Does not Imply Integrability, Report no. WS-482, Free University of Amsterdam (1997). C. Eichler-Liebenow, P.J. van der Houwen, B.P. Sommeijer, Analysis of approximate factorization in iteration methods, CWI Report MAS-R9718 (1997). C.T.H. Everaars and B. Koren, Using coordination to parallelize sparsegrid methods for 3D CFD problems, CWI Report SEN-R9705 (1997). H.C. Hemker, P.W. Hemker and W.J.H. Stortelder Mathematical modelling in blood coagulation, Preprint, CWI Report MAS-R9720 (1997). P.W. Hemker and B. Koren, Standard multigrid techniques for CFD, 28th Computational Fluid Dynamics, VKI LS 1997-02, Von Karman Institute for Fluid Dynamics, Rhode-Saint-Genèse (1997). P.W. Hemker, B. Koren, W.M. Lioen, M. Nool and H.T.M. van der Maarel Solution-adaptive multigrid for steady gas dynamics problems, 28th Computational Fluid Dynamics, VKI LS 1997-02, Von Karman Institute for Fluid Dynamics, Rhode-Saint-Genèse (1997). K.J. in ’t Hout, Ch. Lubich Periodic orbits of delay differential equations under discretizations, Report no. TW-97-02, Department of Mathematics, Leiden University (1997). P.J. van der Houwen and B.P. Sommeijer, The use of approximate factorization in stiff ODE solvers, CWI Report MAS-R9732 (1997). W. Hundsdorfer, Stability of approximate factorization with θ-meth ods, CWI Report MAS-R9721 (1997). M.A. Kirkilionis, O. Diekmann, B. Lisser, M. Nool, A.M. de Roos, B.P. Sommeijer Numerical continuation of equilibria of physiologically structured population models.I. Theory, CWI Report MAS-R9714 (1997). B. Koren and P.W. Hemker, Godunov-type schemes, sparse grids and 3D CFD, Godunov’s Method for Gas Dynamics: Current Applications and Future Developments, Department of Aerospace Engineering, The University of Michigan, Ann Arbor, MI (1997). B. Koren and A.C.J. Venis, A level-set method for moving material-void interfaces, CWI Report MAS-R9731 (1997).

W. Loeve, J. Verwer, E. Snijdoodt, A. ten Dam, Classical HPCN geared to application in industry, Report NLR TP 97229 L, National Aerospace Laboratory, The Netherlands (1997). J. Noordmans and P.W. Hemker, Convergence results for 3D sparse grid approaches, CWI Report MAS-R9713 (1997). J.D. Pinter, W.J.H. Stortelder and J.J.B. de Swart, Computation of elliptic Fekete point sets, CWI Report MAS-R9705 (1997). J.A. Sanders, J.P. Wang, On the Integrability of Polynomial Scalar Evolution Equations, Report no. WS-467, Free University of Amsterdam (1997). J.A. Sanders, J.P. Wang, On heriditary recursion operators, Report no. WS-445, Free University of Amsterdam (1997). B.P. Sommeijer, L.F. Shampine, J.G. Verwer, RKC: an explicit solver for parabolic PDEs, CWI Report MAS-R9715 (1997). E.J. Spee, A.C. Petersen, H. van Dop, W. Hundsdorfer, Sensitivity of Atmospheric Transport Model Performance to Numerical Advection Schemes and Resolution, CWI Report MAS-R9710 (1997). E.J. Spee, J.G. Verwer, P.M. de Zeeuw, J.G. Blom, W. Hundsdorfer, A Numerical Study for Global Atmospheric Transport-Chemistry Problems, CWI Report MAS-R9702 (1997). F. Straetemans, Resolvent conditions for discretizations of time dependent diffusion-convection-reaction equations, Report no. TW-97-05, Dept. of Mathematics, Leiden University (1997). J.J.B. de Swart and G. Söderlind, On the construction of error estimates for implicit Runge-Kutta methods, CWI Report MAS-R9704 (1997). S. Tracogna, B. Welfert, Numerical differentiation and Peano kernel functions, Report TW-97-01, Department of Mathematics, Leiden University (1997). S. Tracogna, B. Welfert Two-step Runge–Kutta methods: theory and practice, Report TW-97-03, Department of Mathematics, Leiden University (1997). S. Tracogna, B. Welfert, Spectral analysis of Generalized Top to Random Shuffles, Report TW-97-04, Department of Mathematics, Leiden University (1997). J.G. Verwer, E.J. Spee, J.G. Blom, W. Hundsdorfer, A second order Rosenbrock method applied to photochemical dispersion problems, CWI Report MAS-R9717 (1997).

Other publications

H. Deconinck and B. Koren, Euler and Navier-Stokes Solvers Using Multi-Dimensional Upwind Schemes and Multigrid Acceleration, European Community Research in Aeronautics, Notes on Numerical Fluid Mechanics, 57, Vieweg, Braunschweig (1997). E.F. Ernst and G.C. Herman, A generalized traveltime method for estimating a laterally varying dispersive velocity field, Book of abstracts, 1 page, Newport RI, 1997, NUWC. P.W. Hemker, Guest editor for a Special Issue on Multilevel Methods of Applied Numerical Mathematics, Volume 23, Number 1, February 1997. (ISSN 0168-9274). P.W. Hemker and B. Koren, Standard multigrid techniques for CFD, 28th Computational Fluid Dynamics, VKI LS 1997-02, Von Karman Institute for Fluid Dynamics, Rhode-Saint-Genèse (1997). P.W. Hemker, B. Koren, W.M. Lioen, M. Nool and H.T.M. van der Maarel, Solution-adaptive multigrid for steady gas dynamics problems, 28th Computational Fluid Dynamics, VKI LS 1997-02, Von Karman Institute for Fluid Dynamics, Rhode-Saint-Genèse (1997) P.W. Hemker, B. Koren and J. Noordmans, 3D multigrid on partially ordered sets of grids, 28th Computational Fluid Dynamics, VKI LS 1997-02, Von Karman Institute for Fluid Dynamics, Rhode-Saint-Genèse (1997). P.W. Hemker, G.I.Shishkin, L.P.Shishkina, On discrete methods arising in the numerical solution of singularly perturbed convection-diffus ion problems, Abstracts of the International Algebraic Conference dedicated to the memory of D.K. Faddeev, St. Petersburg, June 24-30, 1997, p. 56. P.W.Hemker, G.I.Shishkin, L.P.Shishkina, The use of defect correction on unified time grids for the solution of parabolic singular perturbation problems, Abstracts of the Sixth International Colloquium on Numerical Analysis and Computer Science with Applications, Plovdiv, August 13-17, 1997, p. 54. P.W.Hemker, G.I.Shishkin, L.P.Shishkina, A numerical method on floating meshes for singularly perturbed problems with a concentrated disturbance of the initial data, Abstracts of the International Workshop on the Numerical Solution of Thin Layer Phenomena, Amsterdam, November 20-21, 1997, p. 20. P.W.Hemker, G.I.Shishkin, L.P.Shishkina, ϵ-uniform schemes with highorder time-accuracy for parabolic singular perturbation problems. The Neumann problem, submitted for publication in Proceedings of the Sixth International Colloquium on Numerical Analysis and Computer Science with Applications (1997). B. Koren and P.W. Hemker, Godunov-type schemes, sparse grids and 3D CFD, Godunov’s Method for Gas Dynamics: Current Applications and Future Developments, Department of Aerospace Engineering, The University of Michigan, Ann Arbor, MI (1997) W.M. Lioen, J.J.B. de Swart and W.A. van Veen, Test set for IVP solvers,

available on WWW at URL: http://www.cwi.nl/cwi/projects/IVPtestset/.

W.M. Lioen, J.J.B. de Swart and W.A. van Veen, PSIDE users’ guide,

available on WWW at URL: http://www.cwi.nl/cwi/projects/PSIDE/.

J.A. Sanders, J.P. Wang, On the (non)existence of conservation laws for evolution equations, Booktitle: The First Electronics Symposium of the Chinese Scholars in the Netherlands, Eindhoven, (1997), 103–106. J.J.B. de Swart, W.M. Lioen and W.A. van der Veen, Specification of PSIDE,

available on WWW at URL: http://www.cwi.nl/cwi/projects/PSIDE/.

K. Wapenaar, E. Verschuur, A. Koek and F. Ernst, Transformation of marine data to pseudo land data, A.J. Berkhout (ed.), in: Delphi: The Acquisition and Preprocessing Programme. Vol. I, Delft 1997, p. 14.1-14.10.

Stochastics

3.1. Probability

Papers in journals and proceedings

P.C. Allaart, An invariant-sum characterization of Benford’s Law; Journal of Applied Probability, 34, 1997, pp. 288-291. A.A. Balkema, Stopping times and sample path regularity, J. Math. Sci. 83, (1997), 352–355. T. Bedford, On the use of minimally informative copulae in competing risk problems, in Proceedings of Mathematical methods in reliability 16-19 September 1997, RENEL Energy Info-Documentation Centre CIDE, Bucharest, 1997.

T. Bedford, A. Fisher, Ratio geometry, rigidity, and the scenery process for hyperbolic Cantor sets Ergod. Th. & Dynam. Syst., 17, 531-564, 1997. T. Bedford, A.M.H. Meeuwissen, Minimally informative distributions with given rank correlation for use in uncertainty analysis, J. of Statist. Comput. and Simul., 57,1997, 143-174. T. Bedford, I. Meilijson, A characterisation of marginal distributions of (possibly dependent) stochastic variables which right-censor each other. Annals of Statistics, 25, 1622-1645, 1997. J. van de Berg, A constructive mixing condition for 2-d Gibbs measures with ramdom interactions, The Annals of Probability 25, 1316-1333 (1997). J. van de Berg, T. Bedford, A remark on the Van Lieshout and Baddeley J-function for point processes, Adv.Appl.Prob. 29, 19-25 (1997). J. van de Berg, L.E. Meester and R. Prasad, Cochannel interference under Raleigh fading and log-normal shadowing. Proceedings IEEE fifth symposium on Communications and Vhicular Technology in the Benelux, 1997,p. 77-81. J. van de Berg, R. Meester and D. White, Dynamic Boolean models Stochastic Processes and their Applications 69, 247-257 (1997). I.S. Cheng, L. de Haan and X. Huang, Rate of convergence of intermediate order statistics. J. Theoretical Probability 10, (1997) 1-23. R.M. Cooke, “Uncertainty modeling: examples and issues”, Safety Science Vol. 26, no. 12, pp. 49-60, 1997. R.M. Cooke & J. Paulsen, “Concepts for Measuring Maintenance Performance and Methods for Analysing Competing Failure Modes”, Reliability Engineering and System Safety, Vol. 55, no. 2, p. 135-141, 1997. Selected from Proceedings ESREL 94, 9th International Conference Reliability and Maintainability 30/5/94. F.M. Dekking, What is the long range order in the Kolakoski sequence? The Mathematics of Long-Range Aperiodic Order. (Waterloo, Can., 1995.) NATO Adv. Sci. Inst. (Eds.: R.V.Moody and J.Patera), Kluwer Acad. Publ., Dordrecht, 1997, 115-125. F.M. Dekking, R.M. Burton, T. Coffey and K. Hyman, Fractal percolation with neighbour interaction. In: J.Levy-Vehel, E.Lutton et al., Fractals in Engineering, From theory to Industrial Applications, Springer, London, 1997, 106-114. F.M. Dekking, J.J. Kalker and E.A.H. Vollebregt, Simulation of rough elastic contacts. J. Appl. Mechanics 64, 1997, 361-368.

F.M. Dekking, E.R. Speer, On the shape of the wavefront of branching random walk . In: Classical and Modern Branching Processes, (Eds. K.B.Athreya and P.Jagers), IMA Volumes in Mathematics and its Applications Vol.84, Springer, New York, 1997, 73-88. J.H.J. Einmahl and L. de Haan, Estimating the spectral measure of an extreme value distribution Stochastic Processes and their Applications 70 (1997) 143-171 J.H.J. Einmahl, L. de Haan and A. Sinha, Estimation of the spectral measure of an extreme-value distribution. Stoch. Prod. Appl. 70, (1997) 143-171. A. Ermakov, Exact probabilities and asymptotics for the one-dimensional coalescing ideal gas, Stochastic Processes and their Applications v.71 (1997), pp.275-284 J.L. Geluk, Tails of subordinated laws: the regularly varying case, Stoch. Proc. Appl., 61, 147-161, (1996). J.L. Geluk, On the domain of attraction of e^{−e^{−x}}, Stat. Prob. Letters, 31, 91-95, (1996). J.L. Geluk, A renewal theorem in the finite-mean case, Proc. Amer. Math. Soc., 125, 3407-3413, (1997). J. Geluk, L.de Haan, S. Resnick and C. Starica, Second order regular variation, convolution and the central limit theorem. Stoch. Proc. Appl. 69, (1997) 139-159. L.H.J. Goossens, R.M. Cooke, “Applications of some risk assessment techniques: Formal expert judgement and accident sequence precursors”, Safety Science Vol. 26, no. 1/2, pp. 35-47, 1997. L.H.J. Goossens, R.M. Cooke, B.C.P. Kraan, F.T. Harper, M. Young, L. Miller, S.G. Hora and J. Päsler-Sauer, ”Uncertainty analysis of nuclear accident consequence codes: A formal expert judgement approach”, In: O. Renn (Ed.), Risk analysis and management in a global economy, Vol. 1.1. Risk Management in Europe: New challenges for an industrial world, SRA-Europa, Stuutgart 1997, Vol. 1.1, pp. 779-793. L. de Haan and L. Peng, Rate of convergence for bivariate extremes. J. Multivariate Analysis 61, (1997) 195-230. L. de Haan and L. Peng, Slow convergence to normality: an Edgeworth expansion without third moment. Prob. and Math Stat.,17, (1997) 395-406. K. van Harn, F.W. Steutel, A generalization of the life-time pairs of renewal theory; Chapter 5 in: Advances in the Theory and Practice of Statistics: A volume in honour of Samuel Kotz, John Wiley & Sons, 1997, pp. 89-105. M. Harris, M.S. Keane, Random coin tossing, Probability Theory and Related Fields 109, (1997) 27–37. G. Hooghiemstra, P.E. Greenwood, The domain of attraction of the α-sun operator for type II and type III distributions. Bernoulli 1997 3, 479-489. G. Hooghiemstra, M.S. Keane, A central limit theorem for sums of correlated products, Statistica Neerlandica 51, (1997)23–34. G. Hooghiemstra, L. Meester, The extremal index in ten seconds. J. Appl. Probab. 1997 Vol 34 3, 818-822. G. Hooghiemstra, L.E. Meester and R. Prasad, Cochannel interference under Raleigh fading and log-normal shadowing. Proceedings IEEE fifth symposium on Communications and Vehicular Technology in the Benelux, 1997, p. 77-81. J.L. Mijnheer, Asymptotic inference for AR(1) processes with (non-normal) stable errors I. J. Math. Sciences 83 (1997), pp 401-406. J.L. Mijnheer, Asymptotic inference for AR(1) processes with (non-normal) stable errors II - Properties of some stochastic integrals. Theory of Stochastic Processes 2 (18) (1996), pp 201-205. J.L. Mijnheer, Asymptotic inference for AR(1) processes with (non-normal) stable errors III - Stable Ornstein-Uhlenbeck processes and integrals of stable O-U processes. Stochastic Models 13 (1997), pp 661-672. A.G. Pakes and F.W. Steutel, On the number of records near the maximum, Austral. J. Statist. 39(2), 1997, 179-192. A. Sinha, J.H.J. Einmahl and L. de Haan, Estimating the spectral measure of an extreme value distribution Stochastic Processes and their Applications 70 (1997) 143-171 F.W. Steutel and M.J.A. van Eenige, Note on the approximation of distributions on ZZ_{+} by mixtures of negative binomial distributions, COMM. STATIST. Stochastic Models, 13(2),1997, 271-274.

Ph.D. Theses

A. Ermakov, Percolation and coalescing particle systems, Delft University of Technology, 1997. M. Harris Coin Tossing, Random Mirrors and Dependent Percolation: Three Paradigms of Phase Transition, Delft University of Technology, 1997.

A. Sinha, Estimating Failure Probability when failure is rare: multidimensional case, Erasmus University Rotterdam 1997.

Reports and pre-prints

P. C. Allaart, Minimax risk inequalities for the location-parameter classification problem. WS-475, Free University Amsterdam, 1997. P.C. Allaart, G. Grimmett and R. Schinazi, Dependent random graphs and spatial epidemics, preprint. J. Bertoin, K. van Harn, F.W. Steutel, Renewal theory and level passage by subordinators, Memorandum COSOR 97–15, Faculteit Wiskunde en Informatica, Technische Universiteit Eindhoven, juni 1997, 5 pages, preprint. J. Brown, L.H.J. Goossens, B.C.P. Kraan, R.M. Cooke, J.A. Jones, J. Randall, F.T. Harper, F.E. Haskin, M.L. Abbott, M.L. Young, S.C. Hora and A. Rood, Probabilistic Accident Consequence Uncertainty Analysis, NUREG/CR-6523 EUR 16771, Vol. 1. Main Report June 1997. Vol. 2 Appendices. Geluk and L. de Haan, Stable probability distributions and their domains of attraction: a direct approach, (1997) preprint. L.H.J. Goossens, J. Boardman, F.T. Harper, B.C.P. Kraan, R.M. Cooke, M.L. Young, J.A. Jones and S.C. Hora, Probabilistic Accident Consequence Uncertainty Study: Uncertainty Assessment For Deposited Material And External Doses. Prepared for U.S. Nuclear Regulatory Commission and Commission of European Communities, NUREG/CR-6526 EUR 16772 SAND97-2323 Washington/USA, and Brussels-Luxembourg, 1997, published 1997, Vol. 1. Main Report. Vol. 2. Appendices. P. Groeneboom, G. Hooghiemstra and H.P. Lopuhaä, The asymptotic normality of the L_{1} error of the Grenander estimator, joint with Report 97-01. L. de Haan, L. Peng and H. Iglesias Pereira, Approximation by penultimate stable laws, (1997) preprint. L. de Haan, L. Peng and T. Themido Pereiro, A bootstrap based method to achieve optimality I estimating the extreme-value index, (1997) preprint. L. de Haan and A.K. Sinha, Estimating the probability of a rare event, (1997) preprint.

K. van Harn and F.W. Steutel, Stationarity of delayed Lévy processes, Memorandum COSOR 97-16, 5 pages. F.E. Haskin, L.H.J. Goossens, F.T. Harper, B.C.P. Kraan and R.M. Cooke, Probabilistic Accident Consequence Uncertainty Study: Early Health Effects Uncertainty Assessment. Prepared for U.S. Nuclear Regulatory Commission and Commission of European Communities, NUREG/CR-6545 EUR 16775 SAND97-2689 Washinton/USA, and Brussels-Luxembourg, 1997, published 1997, Vol. 1. Main Report. Vol. 2. Appendices. G. Hoogiemstra, G. Koole, On the convergence of the power series algorithm, joint report Vrije Universiteit Amsterdam, G. Hooghiemstra H. Lopuhaä, On the maximum of Groeneboom’s local limit process, Report 97-52. M.S. Keane, R.M. Cooke, Considerations on building a generic tool for MAUT, TU Delft report 1997, preprint. M.S. Keane, T. Kamae, A simple proof of the ratio ergodic theorem, (1997) to appear in: Osaka Journal of Mathematics. M.S. Keane, R. Murray, L.-S. Young, Computing invariant measures for expanding circle maps, (1997) to appear in: Nonlinearity. M.S. Keane, J. Serafin, A new proof of the countable generator theorem, (1997) to appear in: Fundamenta Mathematica. J.L. Mijnheer, Asymptotic inference for AR(1) processes with (non-normal) stable errors IV - A note on the case of a negative unit root, RUL technical report TW-96-14.

Other publications

R.M. Cooke, M. Mendel, H. Vrijling (eds.) Engineering Probabilistic Design and Maintenance for Flood Protection, Kluwer Academic Publishers, 1997. R.M. Cooke, J. Dorrepaal and T. Bedford, “Mathematical review of Swedish Bayesian methodology for nuclear plant reliability data bases”, Book: The Practice of Bayesian Analysis , edited by S. French and J.Q. Smith, pp. 25-54, 1997, ISBN 0 340 66240 9, 1997. R.M. Cooke (ed) Technical Committee Uncertainty Modeling: Benchmark workshop, European safety reliability association, Delft 1997, workshop proceedings.

F.M. Dekking, Iterated Function System, In: M.Hazewinkel; Encyclopedia of Mathematics, Suppl.Vol.I, Kluwer Academic Publishers, Dordrecht/Boston/London, 1997, 321. L.H.J. Goossens and R.M. Cooke, “Evaluation of calibration variables for optimising expert judgements in risk assessments”, Proceedings Annual Meeting of the Society for Risk Analysis-Europe, New Risk Frontiers Stockholm 1997. Ed. B.-M. Droftz-Sjöberg, ISSN 1101-9689. Louis Goossens and Roger Cooke, ”Expert judgement and hazard assessments” MEG Bulletin Issue 2, March 1997, Brussels, 3 p. K. van Harn, F.W. Steutel, A generalization of the life-time pairs of renewal theory; Chapter 5 in: Advances in the Theory and Practice of Statistics: A volume in honour of Samuel Kotz. John Wiley & Sons, 1997, p. 89– 105. J.M. van Noortwijk, M. Kok and R.M. Cooke, “Optimal decisions that reduce flood damage along the Meuse: an uncertainty analysis” Book: The Practice of Bayesian Analysis , edited by S. French and J.Q. Smith. pp. 151-172, 1997. ISBN 0 340 66240 9, 1997. J.M. van Noortwijk, M. Kok and R.M. Cooke, ”Optimal maintenance deiesierus for the sea-bed protection of the easten-scheldt barieen”, in Egineering Probabilistic design and mainterance for flood protection, edited by Cooke, R., Mendel, M. and Vryling, K. (eds). pp. 25-57, Kluwer Dordrecht, 1997, ISBN 0-7923-4399-9.

3.2. Statistics

Papers in journals and proceedings

V. Bentkus, F. Götze and W.R. van Zwet, An Edgeworth expansion for symmetric statistics, Annals of Statistics 25, 851–896 1997 E.E.M. van Berkum, H.N. Linssen, D.A. Overdijk, The reduction and ancillary classification of observations, Journal of Statistical Planning and Inference Volume 45 ,1995, pp 347-372. E.E.M. van Berkum, H.N. Linssen, D.A. Overdijk, Inference rules and inferential distribution, Journal of Statistical Planning and Inference Volume 49 ,1996, pp 305-317. P.J. Bickel, F. Götze and W.R. van Zwet, Resampling fewer than n observations: Gains, losses and remedies for losses, Statistica Sinica 7, 1–31 1997 A. di Bucchianico, Representations of stochastic processes, Systems & Control Letters 25 (1995), 343-350. A. di Bucchianico, A. Bottreau, D.E. Loeb, Implementation of an Umbral Calculus Package, MapleTech 2 (1995), 37-41. A. di Bucchianico, D.E. Loeb, A selected survey of umbral calculus, Dynamical Survey 3, Electronic Journal of Combinatorics, 1995. A. di Bucchianico, D.E. Loeb, Sequences of binomial type with persistent roots, J. Math. Anal. Appl. 199 (1996), 39-58. A. di Bucchianico, D.E. Loeb, Operator expansion in the derivative and multiplication by x, Integral Transf. Spec. Func. 4 (1996), 49-68. S. Chakraborti, P. van der Laan, Precedence tests for complete data: an overview and some results, The Statistician (1996) 45, No. 3, pp. 351 - 369. F.P.A. Coolen, P. van der Laan, On indifference zone selection with a preference threshold, Proceedings 50th Session of the International Statistical Institute, Beijing, China, 1995, pp. 1-2. F.P.A. Coolen, P. van der Laan, Statistical selection: Main approaches and a modification with a preference threshold, Proceedings International Symposium on Statistics in Agriculture and Enviromental Research, CIAT, Colombia, 1995, pp. 231-243. R.J.M.M. Does, Industrial statistics at the University of Amsterdam, in Proceedings of the Workshop on Statistics at Universities: Its Impact for Society (1997), A. Zempléni (ed.), Eötvös University Press, Budapest, 57– 64. R.J.M.M. Does, W.A.J. Schippers and A. Trip, A framework for implementation of statistical process control, International Journal of Quality Science 2 (1997), 181–198. F.C. Drost and C.A.J. Klaassen, Efficient estimation in semiparametric GARCH models, J. Econometrics 81 (1997), 193–221. F.C. Drost, C.A.J. Klaassen and B.J.M. Werker, Adaptive estimation in time series models, Ann. Statist. 25 (1997), 786–817. J.H.J. Einmahl, A Bahadur-Kiefer theorem beyond the largest observation, (1995) Journal of Multivariate Analysis 55, 29-38. J.H.J. Einmahl, Extension to higher dimensions of the Jaeschke-Eicker result on the standardized empirical process, (1996) Communications in Statistics - Theory and Methods 25, 813-822.

J.H.J. Einmahl, A short and elementary proof of the main Bahadur-Kiefer theorem, (1996) Annals of Probability 24, 526-531. J.H.J. Einmahl, Poisson and Gaussian approximation of weighted local empirical processes, Stochastic Processes and their Applications, Volume 70, 1997, pp. 31-58. J.H.J. Einmahl, J. Beirlant, Asymptotic confidence intervals for the length of the short t under random censoring, (1995) Statistica Neerlandica 49, 1-8. J.H.J. Einmahl, J. Beirlant, Maximal type test statistics based on conditional processes, (1996) Journal of Statistical Planning and Inference 53, 1-19. J.H.J. Einmahl, P. Deheuvels, On the strong limiting behavior of local functionals of empirical processes based upon censored data, (1996) Annals of Probability 24, 504-525. J.H.J. Einmahl, L. de Haan, A.K. Sinha, Estimating the spectral measure of an extreme value distribution, (1997) Stochastic Processes and their Applications 70, 143-171. J.H.J. Einmahl, F.H. Ruymgaart, Tail processes under heavy random censorship with applications, (1995) Journal of Nonparametric Statistics 5, 223-235. A.J. van Es, A note on the integrated squared error of a kernel density estimator in non–smooth cases, Statistics and Probability Letters, 35 (1997) 241–250. A.J. van Es and A.J. Hoogstrate, How much do plug–in bandwidth selectors adapt to non–smoothness ?, Nonparametric Statistics, 8 (1997) 185–197. S. van de Geer and M. Wegkamp, Consistency of the least squares estimator in non-parametric regression, Annals of Statistics 24, 2513–2523 1997 D. Gilat, R. Helmers, On strong laws for generalized L-statistics with dependent data, Comment. Math.Univ.Carolinae 38,1, 187-192. 1997 R. Helmers, A note on bootstrapping the local time of the empirical process, Statistics and Decisions 15, 295-300. 1997 E.R. van den Heuvel and C.A.J. Klaassen, Spread, estimators and nuisance parameters, Bernoulli 3, (1997) 323–328.

E.R. van den Heuvel, C.G.M. Oudshoorn and R.J.M.M. Does, Six sigma quality programs, Proceedings of the Workshop on Statistics at Universities: Its Impact for Society (1997), A. Zempléni (ed.), Eötvös University Press, Budapest, 83–92. G. Jongbloed, L.M.M. Tijskens, E.C. Wilkinson, Nonlinear models with random coefficients: an example with the logistic growth curve, Proc. of the 12th International Workshop on Statistical Modelling Biel/Bienne, Zwitserland, pp. 294-298. S. Kasper, M. Zivkov, K.C.B. Roes and A.G. Pols, Pharmacological treatment of severely depressed patients: a meta-analysis comparing efficacy ofmirtazapine and amitriptyline, European Neuropsychopharmacology 7 (1997) 115–124. C.A.J. Klaassen and J.A. Wellner, Efficient estimation in the bivariate copula model: normal margins are least favorable, Bernoulli 3 (1997), 55–77. P. van der Laan, Subset selection from exponential populations: determination of the selection constant, Biometrical Journal 37, 1995, pp. 515-521. P. van der Laan, Distributional and efficiency results for subset selection, Journal of Statistical Planning and Inference 54 (1996) 159 - 174. P. van der Laan, Basic results for subset selection, Bulletin International Statistical Institute, ISI Session 51, Tome LVII (1997) 311 - 312. P. van der Laan and C. van Eeden, On using a loss function in selecting the best of two Gamma populations in terms of their scale parameters, Statistics 28 (1996) pp. 355 - 370. E. Mammen and S. van de Geer, Locally adaptive regression splines, Annals of Statistics 25, 387–413 1997 E. Mammen and S. van de Geer, Penalized quasi-likelihood estimation in partial linear models, Annals of Statistics 25, 1014–1035 1997 S.A. Murphy, A.W. van der Vaart, MLE in the proportional odds model, Journal of the American Statistical Association 92, 1997, pp. 968-986. S.A. Murphy, A.W. van der Vaart, Semiparametric likelihood ratio tests, Annals of Statistics, 25, 1997, pp. 1471-1509. H. Putter and W.R. van Zwet, On a set of the first category, Festschrift for Lucien Le Cam, D. Pollard, E. Torgersen and G.L. Yang, editors, Springer Verlag, New York (1997), 315–323.

K.C.B. Roes, Statistical Thinking in Service Processes, in Proceedings of the International Quality Conference, European Quality Week in Hungary, 1997 K.C.B. Roes and D. Dorr, Implementation of Statistical Process Control in service processes, International Journal of Quality Science, 2 (1997), 149-166. W.A.J. Schippers and R.J.M.M. Does, Implementing statistical process control in industry: the role of statistics and statisticians, Proceedings of the Workshop on Statistics at Universities: Its Impact for Society (1997), A. Zempléni (ed.), Eötvös University Press, Budapest, 65–82. A.W. van der Vaart, Superefficiency, Festschrift for L.M. Le Cam, Springer Verlag, New York 1997, pp. 397-410. J.H.M. Verheijen, F.P.A. Coolen and P. van der Laan, Combining two classical approaches for statistical selection, Commun. Statist.- Theory and Methods 26 (1997), 1291 - 1312.

Ph.D. Theses

M. Fiocco, Statistical Estimation for the Supercritical Contact Process, Leiden University, 1997. R.B. Geskus, Estimation of Smooth Functionals with Interval Censored Data, Delft University of Technology, 1997.

Reports and pre-prints

E.E.M. van Berkum, H.N. Linssen, D.A. Overdijk, ”Sampling the reference set” revisited, Memorandum COSOR 96-26, 32 pp. 1996. E.E.M. van Berkum, P.M. Upperman, Some new designs for quantitative factors, Memorandum COSOR 95-03, 28 pp. 1995. P.C. Boon and K.C.B. Roes, Design and analysis issues for crossover designs in phase I clinical trials, 1997, to appear in Journal of Biopharmaceutical Statistics. A. di Bucchianico, M. van de Wiel and P. van der Laan, Symbolic computation and exact distributions of nonparametric test statistics, COSOR memorandum 97-17, Department of Mathematics, Eindhoven University of Technology. S. Chakraborti, P. van der Laan, An overview of precedence-type tests for censored data, to appear in: Biometrical J.

S. Chakraborti, P. van der Laan, Precedence tests for right-censored data: An overview and some results, Memorandum COSOR 96-01, Department of Mathematics and Computing Science, Eindhoven University of Technology. F.P.A. Coolen, P. van der Laan, On indifference zone selection with a preference threshold, Memorandum COSOR 95-02, Eindhoven University of Technology R.J.M.M. Does and A. Zempléni, Establishing a commercial statistical consulting unit at universities, Report of the Department of Mathematics of the Eötvös University, 1997 Budapest. J.H.J. Einmahl, F.H. Ruymgaart, Some results for empirical processes of locally dependent time series, preprint December 1997, 14 blz. J.H.J. Einmahl, M.C.A. van Zuijlen, On the approximation of an integral by a sum of random variables, to appear in Journal of Applied Mathematics and Stochastic Analysis (1998). A.J. van Es, R. Helmers, M. Hušková, On a crossroad of resampling plans: Bootstrapping elementary symmetric polynomials, 15p. CWI report PNA-R9704. A.J. van Es and A. Kok, Simple kernel estimators for certain nonparametric deconvolution problems, Report WINS 97-2, 1997, to appear in Statist. Prob. Lett. M. Fiocco, Statistical estimation for the supercritical contact process, preprint, 1997 S.A. van de Geer, Applications of empirical process theory, Techn. Report TW-97-09 (RUL). R. Helmers, Inference on rare errors using asymptotic expansions and bootstrap calibration, preprint. R. Helmers, R. Zitikis, On estimation of Poisson intensity functions, revised version of Report 270 of the Laboratory for Research in Statistics and Probability, Carleton University, Ottawa, to appear in the Annals of the Institute of Statistical Mathematics. B.Y. Jing, G. Qin, R. Helmers, Saddlepoint approximations to the trimmed mean, preprint. A.J. Koning and R.J.M.M. Does, Cusum charts for preliminary analysis of individual observations, Report 9727/A, 1997, Econometric Institute, Erasmus Universiteit Rotterdam.

M.J. van der Laan and P. van der Laan, Subset selection from generalized logistic populations, Technical Report No. 63, (1997), Group in Biostatistics, University of California, Berkeley. M.J. van der Laan and P. van der Laan, Subset selection from generalized logistic populations, Memorandum COSOR 97-12 ( 18 pagina’s), Department of Mathematics and Computing Science, Eindhoven University of Technology. P. van der Laan, Statistical Selection: A way of thinking, Memorandum COSOR 95-25, Eindhoven University of Technology. P. van der Laan, Distributional and efficiency results for subset selection, Memorandum COSOR 96-07, Revised version of 94-03. Department of Mathematics and Computing Science, Eindhoven University of Technology. P. van der Laan, F.P.A. Coolen, An overview of a generalization in Statistical Selection, Memorandum COSOR 95-20, Eindhoven University of Technology. P. van der Laan and C. van Eeden, On selecting the best of two normal populations using a loss function, Memorandum COSOR 95-14, Eindhoven University of Technology. P. van der Laan and C. van Eeden, On using a population in selecting the best of two Gamma populations in terms of their scale parameters, Memorandum COSOR 95-26 Eindhoven University of Technology, Technical Report 153, Dept. of Statistics, University of British Columbia, Vancouver, Canada, 1995. P. van der Laan and C. van Eeden, On using a loss function in selecting the best of two Gamma populations in terms of their scale parameters, To appear in Biometrical Journal. P. van der Laan and C. van Eeden, Subset selection for logistic populations using loss functions: Some basic probabilities, Memorandum COSOR 97-14 ( 13 pagina’s). Department of Mathematics and Computing Science, Eindhoven University of Technology. P. van der Laan and C. van Eeden, On selecting the best of two normal populations using a loss function, Memorandum COSOR 97-13 ( 18 pagina’s), Department of Mathematics and Computing Science, Eindhoven University of Technology. E.G. Luebeck and M.C.M. de Gunst, A stereological method for the analysis of cell counts in tissue sections using 3-dimensional cellular automata, Rapport WS-470, Faculteit der Wiskunde en Informatica, Vrije Universiteit, Amsterdam, 1997 B. Pauwels, E.E.M. van Berkum, Approximate and exact designs for incomplete quadratic models, Memorandum COSOR 95-04, 14 pp. 1995. B. Pauwels, E.E.M. van Berkum, D-optimal weighted paired comparison designs, Memorandum COSOR 96-30, 15 pp. 1996. H. Putter and W.R. van Zwet, Empirical Edgeworth expansions for symmetric statistics, To appear in: Annals of Statistics. F.S. Radhakishun, J van den Bos, B. van der Heijden, K.C.B. Roes, J.F. Hanlon and the Dutch mirtazapine study group, Daily computer assisted telephone administration of self-rating scale that demonstrate mirtazapine is associated with mild and short-lasting drowsiness and rapid improvement of sleep and mood in depressed patients, 1997, to appear in J. Clin. Psychopharmacology. K.C.B. Roes, R.J.M.M. Does and B.S. Jonker, Effective application of Q(R) charts in low volume manufacturing, Report of the Department CPRO of DLO, 1997, Wageningen, Submitted to Quality and Reliability Engineering International. E.S. Tan, A.W. Ambergen, R.J.M.M. Does and Tj. Imbos, Approximations of normal IRT models for change, Report of the Department of Methodology and Statistics, 1997, University of Maastricht, to appear in the Journal of Educational and Behavioral Statistics. M.A. van der Wiel, A. di Bucchianico and P. van der Laan, Symbolic computation and exact distributions of nonparametric test statistics, Memorandum COSOR 97 - 17 (13 pag.). Department of Mathematics and Computing Science, Eindhoven University of Technology.

Other publications

E.E.M. Berkum, Paired comparison model, Encyclopaedia of Mathematics, Supplement I, Kluwer, Dordrecht, 1997, p. 148. E.E.M. Berkum, Bradley-Terry model, Encyclopaedia of Mathematics, Supplement I, Kluwer, Dordrecht, 1997, p. 399-400. A. di Bucchianico, Probabilistic and Analytic Aspects of the Umbral Calculus, CWI tract 119, 1997, 148 pp. R.J.M.M. Does and K.C.B. Roes, Statistische procesbeheersing in de dienstverlening: een onderzoek naar de toepassing van SPC in de dienstverlening, Sigma 5 (1997), 20–25.

J.H.J. Einmahl, Empirical process, Encyclopaedia of Mathematics, Supplement I, Kluwer, Dordrecht, 1997, (blz. 239-240) J.H.J. Einmahl, R.M. Dudley, Vapnik-Chervonenkis class, Encyclopaedia of Mathematics, Supplement I, Kluwer, Dordrecht, 1997, (blz. 478-479) C.A.J. Klaassen, Bonus-Malus in Kwaliteitscontrole, in Vakantiecursus 1997, Rekenen op het Toeval, CWI Syllabus 43, CWI, Amsterdam, 83–102 P. van der Laan, Statistical Selection: A way of thinking, Liber Amicorum prof. F. Kruseman-Aretz, 1995, pp. 209-214. W.R. van Zwet, Comment on “Improving the relevance of economic statistics”, by Yves Franchet, in Economic Statistics: Accuracy, Timeliness and Relevance, US Department of Commerce, Bureau of Economic Analysis, Washington DC, 23–26.1997

Operations Research

4.1. Optimization Theory

Papers in journals and proceedings

S. Ball, A. Blokhuis, F. Mazzocca, Maximal arcs in Desarguesian planes of odd order do not exist, Combinatorica 17 (1997) 31–41. A. Blokhuis, A.E. Brouwer, Determination of the distance-regular graphs without 3-claws, Discrete Mathematics 163 (1997) 225–227. A. Blokhuis, R. Pelikaan, T. Szönyi, Blocking sets of almost Rédei type, Journal of Combinatorial Theory Series A 78 (1997) 141–150. J. Barzilai, F.A. Lootsma, Power relations and group aggregation in the multiplicative AHP and SMART, Journal of Multi-Criteria Decision Analysis 6 (1997) 155-165. B. Baumeister, D.V. Pasechnik, The universal covers of certain semibiplanes, European Journal of Combinatorics 18 (1997) 491–496. A.E. Brouwer, H.M. Mulder, The vertex connectivity of a (0, 2)-graph equals its degree, Discrete Mathematics 169 (1997) 153 - 155. J. Edmonds, M. Laurent, A. Schrijver, A minor-monotone graph parameter based on oriented matroids, Discrete Mathematics 165/166 (1997) 219–226. T. Fleiner, Covering a symmetric poset by symmetric chains, Combinatorica 13 (1997) 339–344.

A. del Fra, A. Pasini, D.V. Pasechnik, A new family of extended generalized quadrangles, European Journal of Combinatorics, 18 (1997) 155– 169. K. Fukuda, T. Terlaky, Criss-cross methods: A fresh view on pivot algorithms, Mathematical Programming Series B 79 (1997) 369–395. C.E.G. van Gennip, J.A.M. Hulshof, F.A. Lootsma, A Multi-Criteria evaluation of diseases in a study for Public Health planning, Eur. J. of Operational Reseacrh 99, 236-24-, 1997 M. de Graaf, A. Schrijver, Making curves minimally crossing by Reidemeister moves, Journal of Combinatorial Theory, Series B 70 (1997) 134– 156. M. de Graaf, A. Schrijver, Decomposition of graphs on surfaces, Journal of Combinatorial Theory, Series B 70 (1997) 157–165. B. He, E. de Klerk, C. Roos, T. Terlaky, Method of approximate centers for semi-definite programming, Optimization Methods and Software 7 (1997) 291–309. R.C. van den Honert, F.A. Lootsma, Group preference aggregation in the multiplicative AHP: the model of the group decision process and Pareto Optimality, European Journal of Operations Research 96 (1996) 363-370. B. Jansen, J.J. de Jong, C. Roos, T. Terlaky, Sensitivity analysis in linear programming: just be careful!, European Journal of Operations Research 101 (1997) 15–28. B. Jansen, C. Roos, T. Terlaky, A Family of Polynomial Affine Scaling Algorithms for Positive Semi–Definite Linear Complementarity Problems, SIAM Journal on Optimization 7 (1997) 126–140. B. Jansen, C. Roos, T. Terlaky, Y. Ye, Improved Complexity Using Higher-Order Correctors for Primal–Dual Dikin Affine Scaling, Mathematical Programming 76 (1997) 117–130. B. Jansen, C. Roos, T. Terlaky, A. Yoshise, Polynomiality of Primal– Dual Affine Scaling Algorithms for Nonlinear Complementarity Problems, Mathematical Programming 78 (1997) 315–345. E. de Klerk, T. Illés, A. de Jong, C. Roos, T. Terlaky, J. Valkó, J.E. Hoogenboom, Optimization of the Reloading Pattern in a Nuclear Reactor, Annals of Operations Research 69 (1997) 65–84. E. de Klerk, C. Roos, T. Terlaky, Initialization in semidefinite programming via a self–dual skew-symmetric embedding, Operations Research Letters 20 (1997) 213–221.

E. de Klerk, C. Roos, T. Terlaky, Polynomial primal-dual affinescaling algortihms in semidefinite programming, Journal on Combinatorial Optimization 2 (1997) 47–68. M. Laurent, Max-cut problem, In: M. Dell’Amico, F. Maffioli and S. Martello (editors), Annotated Bibliographies in Combinatorial Optimization, John Wiley, pages 241–259. M. Laurent, The real positive semidefinite completion problem for seriesparallel graphs, Linear Algebra and its Applications 252 (1997) 347–366. M. Laurent, Cuts, matrix completions and graph rigidity, Mathematical Programming 79 (1997) 255–283. M. Laurent, S. Poljak, F. Rendl, Connections between semidefinite relaxations of the max-cut and stable set problems, Mathematical Programming 77 (1997) 225–246. F.A. Lootsma, Comment on ”The European school of MCDA: Emergence, basic features and current work” by B. Roy and D. Vanderpooten, Journal of Multi-Criteria Decision Analysis 5, 37-38, 1996. F.A. Lootsma, The decision maker and the analyst in MCDA, Journal of Multi-Criteria Decision Analysis 5, 167-168, 1996. F.A. Lootsma, MCDA, a new technology, Journal of Multi-Criteria Decision Analysis 5, 245-246, 1996. F.A. Lootsma, A model for the relative importance of the criteria in the multiplicative AHP and SMART, Eur. J. of Operational Research 94, 467-476, 1996. F.A. Lootsma, Multi-Criteria decision analysis in a decision tree, Proceedings of the 4th International Symposium on the AHP, Simon Fraser University, Barnaby, B.C. Canada, 1996, pp. 251-261 . F.A. Lootsma, Fuzzy Smart, Proceedings of the 4th European Congress on intelligent techniques, Aachen, Germany, 1996, pp. 1391-1393. F.A. Lootsma, Multi-Criteria decision analysis in a decision tree, Eur. J. of Operational Research 101, 442-451, 1997. F.A. Lootsma, T.W. Athan, P.Y. Palambros, Controlling the search for a compromise solution in multo-objective optimization, Engineering Optimization 25, 65-81, 1995. F.A. Lootsma, H. Schuijt, The multiplicative AHP, SMART and ELEC-TRE in a common context, Journal of Multi-Criteria Decision Analysis 6 (1997) 185-196.

H. van Maaren, T. Terlaky, Inverse barriers and CES-functions in linear programming, Operations Research Letters 20 (1997) 15–20. F.R. McMorris, H.M. Mulder, Subpath acyclic digraphs, Discrete Math. 154 (1996) 189 201. H.M. Mulder, The Majority Strategy on Graphs, Discrete Applied Mathematics 80 (1997) 97 - 195. A. Schrijver, Minor-monotone graph invariants, In: R.A. Bailey (editor), Surveys in Combinatorics Cambridge University Press, Cambridge, pp. 163– 196. J. Warners, T. Terlaky, C. Roos, B. Jansen, A Potential Reduction Approach to the Frequency Assignment Problem, Discrete Applied Mathematics 78 (1997) 252–282. J. Warners, T. Terlaky, C. Roos, B. Jansen, Potential Reduction Algorithms for Structured Combinatorial Optimization Problems, Operations Research Letters 21 (1997) 55–65.

Ph.D. Theses

E. de Klerk, Interior Point Methods for Semidefinite Programming, 1997, Delft University of Technology.

Reports and pre-prints

C.E.G. van Gennip, J.A.M. Hulshof, F.A. Lootsma, A multi-criteria evaluation of diseases in a study for Public-Health planning, TWI Report 96-53, Delft Univ. of Technology 1996, to appear in Eur. J. of Operational Research. W. Imrich, S. Klavzar, H.M. Mulder, Median graphs and triangle-free graphs, Preprint Series, University of Ljubljana, Vol. 35 (1997) 566. R.E. Jamison, H.M. Mulder, Tolerance intersection graphs on subtrees of a binary tree with constant tolerance 3, Report 9752/B, Econometrisch Instituut EUR, 1997, 21 pp. S. Klavzar, H.M. Mulder, Graphs that are locally a hypercube, Preprint Series, University of Ljubljana, Vol. 33 (1995), 471, 14 pp. S. Klavzar, H.M. Mulder, Median graphs: characterizations, location theory and related structures, Report 9641/B, Econometrisch Instituut EUR, 1996, 20 pp.

E. de Klerk, C. Roos, T. Terlaky, Infeasible start semidefinite programming algorithms via self-dual embeddings, Delft University of Technology, Department of Technical Mathematics and Informatics, Report No. 97–10. E. de Klerk, C. Roos, T. Terlaky, Primal–dual potential reduction methods for semidefinite programming using affine–scaling directions, Delft University of Technology, Department of Technical Mathematics and Informatics, Report No. 97–29. F.A. Lootsma, Preferences in perspective, TWI-Report 96-?, Delft University of Technology, 1996. F.A. Lootsma, H. Schuijt, Fairness and equity via concepts of multicriteria decision analysis, TWI-Report 96-101, Delft University of Technology, 1996. Z.-Q. Luo, C. Roos, T. Terlaky, Complexity analysis of a logarithmic barrier decomposition method for semi-infinite linear programming, Delft University of Technology, Department of Technical Mathematics and Informatics, Report No. 97-34. M.A. Morgana, H.M. Mulder, Induced paths with a common vertex, Report 9753/B, Econometrisch Instituut EUR, 1997, 26 pp. F.R. McMorris, H.M. Mulder, R.C. Powers, The median function on median graphs and semilattices, Report 9651/A, Econometrisch Instituut EUR, 1996, 13 pp. F.R. McMorris, H.M. Mulder, F.S. Roberts, The median procedure on graphs, Report 9535/B, Econometrisch Instituut EUR, 1995, 25 pp. H.M. Mulder, The majority strategy on graphs, Report 9640/B, Econometrisch Instituut EUR, 1996, 12 pp. D.V. Pasechnik, C.E. Praeger, On transitive permutation groups with primitive subconstituents, Delft University of Technology, Department of Technical Mathematics and Informatics, Report No. 97-27. J.M. Peng, Derivative-free methods for monotone variational inequality and complementarity problems, Delft University of Technology, Department of Technical Mathematics and Informatics, Report No. 97-56. J.M. Peng, A smoothing function and its applications, Delft University of Technology, Department of Technical Mathematics and Informatics, Report No. 97-57.

A.J. Quist, E. de Klerk, C. Roos, T. Terlaky, Copositive relaxation for general quadratic programming, Delft University of Technology, Department of Technical Mathematics and Informatics, Report No. 97–50. R. Ramanathan, F.A. Lootsma, Fairness issues in group decision making and a model using the multiplicative AHP, TWI-Report 96-05, Delft University of Technology, 1996.

Other publications

A.B. Berkelaar, C. Roos, T. Terlaky, The optimal set and optimal partition approach to linear and quadratic programming, In: Tomas Gal and Harvey J. Greenberg (editors), Advances in Sensitivity Analysis and Parametric Programming, pp. 6.1–6.44. Kluwer Academic Publishers, Boston/Dordrecht/London, 1997. M. Deza, M. Laurent, Geometry of Cuts and Metrics, Springer-Verlag, Berlin, 1997. T. Illés, C. Roos, T. Terlaky, Polynomial Affine-Scaling Algorithms for P_{∗}(κ) Linear Complementarity Problems, In: P. Gritzmann, R. Horst, E. Sachs and R. Tichatschke (editors), Recent Advances in Optimization, pp. 119–137. Proceedings 8th French–German Conference on Optimization (July 21–26, 1996, Trier, Germany). Lecture Notes in Economics and Mathematical Systems 452, Springer-Verlag, Berlin/Heidelberg/New York, 1997. B. Jansen, Interior Point Techniques in Optimization. Complexity, Sensitivity and Algorithms, Kluwer Academic Publishers, Dordrecht, 1997. E. de Klerk, C. Roos, T. Terlaky, A short survey on semidefinite programming, In: W.K. Klein Haneveld, O.J. Vrieze and L.C.M. Kallenberg (editors), Ten years LNMB, pp. 323–339. CWI Tract 122, Centrum voor Wiskunde en Informatica, Amsterdam, The Netherlands, 1997. F.A. Lootsma, Fuzzy logic for planning and decision making, Kluwer Academic Publishers, Dordrecht, 1997. C. Roos and T. Terlaky, Advances in Linear Optimization, Chapter 7 in M. Dell’Amico, F. Maffiolo and S. Martello (editors), Annotated Bibliographies in Combinatorial Optimization, John Wiley & Sons, Ltd, pp. 95–114, 1997. C. Roos, T. Terlaky, J.-Ph. Vial, Theory and Algorithms for Linear Optimization. An Interior Point Approach, John Wiley & Sons, Chichester, 1997.

Papers in journals and proceedings

I.J.B.F. Adan, G.J. van Houtum, J. van der Wal, The symmetric longest queue system. Stochastic Models, Vol. 13, 1997, pp. 105-120. I.J.B.F. Adan en J.A.C. Resing, Steady state probabilities for circular Markov chains. The Mathematical Scientist 22, 1997, pp 27-31. I.J.B.F. Adan, J. Wessels, Shortest expected delay routing for Erlang servers. Queueing Systems, Vol. 23, 1996, pp. 77-105, (appeared in 1997). E. Altman and A. Hordijk, Applications of Borovkov’s renovation theory to non-stationary stochastic recursive sequences and their control, Advances in Applied Probability 29 (1997) 388-413. E. Altman, A. Hordijk and F.M. Spieksma, Contraction conditions for average and discount optimality in countable state Markov games with unbounded rewards, Mathematics of Operations Research 22 (1997) 588-618. S.C. Borst, O.J. Boxma, Polling models with and without switchover times, Operations Research 45, 1997, 536-543. R.J. Boucherie, O.J. Boxma, K. Sigman, A note on negative customers, GI/G/1 workload, and risk processes, Probability in the Engineering and Informational Sciences 11, 1997, 305-311. O.J. Boxma, Regular variation in a multi-source fluid queue, Teletraffic Contributions for the Information Age, (Proc. ITC-15). V. Ramaswami, P.E. Wirth (Eds.), 1997, 391-402. O.J. Boxma, D.G. Down, Dynamic server assignment in a two-queue model, EJOR 103, 1997, 101-115. O.J. Boxma, U. Yechiali, An M/G/1 queue with multiple types of feedback and gated vacations, J. Appl. Probability 34, 1997, 773-784. J.W. Cohen, On the determination of the stationary distribution of a symmetric clocked buffered switch, Teletraffic Contributions for the Information Age (Proc. ITC-15). V. Ramaswami, P.E. Wirth (Eds.), North-Holland Publ. Cy., Amsterdam, 1997, 297-308. J.W. Cohen, D.G. Down, On the role of Rouché’s theorem in queueing analysis, Queueing Systems 23, 1996, 281-291. A.S. Gajrat, A. Hordijk, V.A. Malyshev and F.M. Spieksma, Fluid approximations of Markov decision chains, Markov Processes and Related Fields 3 (1997) 129-150.

M. Haviv and J. van der Wal, Equilibrium strategies for processor sharing and random queues with relative priorities. Probability in the Engineering and Informational Sciences 11 (1997), pp. 403 - 412. A. Hordijk and A. Loeve, Optimal noncyclic server allocation in a polling model, Proceedings of the 36th IEEE Conference on Decision and Control (1997) 2941-2945. A. Hordijk, O. Passchier and F.M. Spieksma, Optimal service control against worst case admission policies: a multichained stochastic game, Mathematical Methods of Operations Research 45 (1997) 281-301. G.J. van Houtum, I.J.B.F. Adan, J. Wessels, W.H.M. Zijm, Performance analysis of parallel machines with a generalized shortest queue arrival mechanism. Proceedings of the International Workshop on Performance Evaluation and Optimization of Production Lines, University of the Aegean, Samos, 1997, pp. 157-172. G.M. Koole, O. Passchier, Optimal control in light traffic Markov decision processes. ZOR - Mathematical Methods of Operations Research 45: 63–79, 1997. G.M. Koole, Assigning a single server to inhomogeneous queues with switching costs. Theoretical Computer Science 182: 203–216, 1997. G.M. Koole, On the use of the power series algorithm for general Markov processes, with an application to a Petri net. INFORMS Journal on Computing 9: 51–56, 1997 V.A. Malyshev and F.M. Spieksma, Dynamics in binary neural networks with a finite number of patterns. Part I: Asynchronous zero temperature dynamics, Math. Physics Journal (1997). S.A.E. Sassen and J. van der Wal, The response-time distribution in a real-time database with optimistic concurrency control and exponential execution times. Proceedings of the 15th International Teletraffic Congress - ITC 15, eds. V. Ramaswami and P.E. Wirth, Washington, DC, USA, 22-27 June, 1997, pp. 145-156. North-Holland. P.D.V. van der Stok, S.A.E. Sassen, M.P. Bodlaender, J. van der Wal and A.T.M. Aerts, Performance modeling of real-time database schedulers. Chapter 15 in: Real-Time Database Systems: Issues and Applications, edited by A. Bestavros, K.J. Lin and S.H. Son, Kluwer Academic Publishers, 1997.

Reports and pre-prints

I.J.B.F. Adan, Explicit solutions for queueing problems, Memorandum COSOR 97-05 (10 pag.) I.J.B.F. Adan, E.A. van Doorn, J.A.C. Resing and W.R.W. Scheinhardt, Analysis of a single-server queue interacting with a fluid reservoir, COSOR memorandum 97-20, October 1997, 23 pag. E. Altman, B. Gaujal and A. Hordijk, Admission control in stochastic event graphs, INRIA research report No. 3179 (1997) (Revision: RUL report TW 97-06). E. Altman, B. Gaujal and A. Hordijk, Multimodularity, convexity and optimization properties, INRIA research report No. 3181 (1997) (Revision: RUL report TW 97-07). E. Altman, B. Gaujal and A. Hordijk, Balanced sequences and optimal routing, INRIA research report No. 3180 (1997) (Revision: RUL report TW 97-08). O.J. Boxma, J.W. Cohen, Heavy-traffic analysis for the GI/G/1 queue with heavy-tailed distributions, PNA-R9710 O.J. Boxma, V. Dumas, Fluid queues with long-tailed activity period distributions, PNA-R9705 O.J. Boxma, V. Dumas, The busy period in the fluid queue, PNA-R9718 J.W. Cohen, On the M/G/1 queue with heavy-tailed service time distributions, PNA-R9702 J.W. Cohen, A heavy-traffic theorem for the GI/G/1 queue with a Paretotype service time distribution, PNA-R9709 J.W. Cohen, The M/G/1 fluid model with heavy-tailed message length distributions, PNA-R9714 J.W. Cohen, Heavy-traffic limit theorems for the heavy-tailed GI/G/1 queue, PNA-R9719 M. Haviv and J. van der Wal, Equilibrium strategies for processor sharing and random queues with relative priorities, Probability in the Engineering and Informational Sciences 11 (1997), pp. 403 - 412. G. Hooghiemstra, G.M. Koole, On the convergence of the power series algorithm. Technical Report WS-487, Vrije Universiteit Amsterdam, 1997. G.M. Koole, Z. Liu, Stochastic bounds for queueing systems with multiple on-off sources. Technical Report WS-471, Vrije Universiteit Amsterdam, 1997.

G.M. Koole, On the static assignment to parallel servers. Technical Report WS-477, Vrije Universiteit Amsterdam, 1997. G.M. Koole, P.D. Sparaggis, D. Towsley, Minimizing response times and queue lengths in systems of parallel queues. Technical Report WS-479, Vrije Universiteit Amsterdam, 1997. G.J. van Houtum, W.H.M. Zijm, I.J.B.F. Adan, and J. Wessels, Bounds for performance characteristics; a systematic approach via cost structures, Memorandum COSOR 97-08 (16 pag.) R. Núñez Queija, Steady-state analysis of a queue with varying service rate, PNA-R9712 S.A.E. Sassen and J. van der Wal, The response time distribution in a real-time database with optimistic concurrency control and exponential execution times, Memorandum COSOR 97-03, EUT, Eindhoven, 1997, pp. 12. S.A.E. Sassen and J. van der Wal, The response time distribution in a real-time database with optimistic concurrency control and constant execution times, Memorandum COSOR 97-07, EUT, Eindhoven, 1997, pp. 23. S.A.E. Sassen and J. van der Wal, The M|G|∞ queue with OCC, Memorandum COSOR 97-18, EUT, Eindhoven, 1997, pp. 19. S.A.E. Sassen and J. van der Wal, Performance analysis of a real-time database with optimistic concurrency control, Memorandum COSOR 97-21, EUT, Eindhoven, 1997, pp. 10. S.A.E. Sassen and J. van der Wal, The response time distribution in a real-time database with optimistic concurrency control and general execution times, Memorandum COSOR 97-22, EUT, Eindhoven, 1997, pp. 20. S.A.E. Sassen and J. van der Wal, The response-time distribution in a real-time database with optimistic concurrency control and exponential execution times, Proceedings of the 15th International Teletraffic Congress - ITC 15, eds. V. Ramaswami and P.E. Wirth, Washington, DC, USA, 22-27 June, 1997, pp. 145-156. North-Holland. P.D.V. van der Stok, S.A.E. Sassen, M.P. Bodlaender, J. van der Wal and A.T.M. Aerts, Performance modeling of real-time database schedulers, Chapter 15 in: Real-Time Database Systems Issues and Applications, edited by A. Bestavros, K.J. Lin and S.H. Son, Kluwer Academic Publishers, 1997.

Other publications I.J.B.F. Adan, Explicit solutions for queueing problems, In: Ten years LNMB, W.K. Klein Haneveld, O.J. Vrieze, L.C.M. Kallenberg (eds.), CWI Tract 122, Mathematical Centre, Amsterdam, 1997, p. 382. A. Hordijk and A. Loeve, Analysis of a customer assignment model with decentralized control, In: Ten years LNMB, W.K. Klein Haneveld, O.J. Vrieze, L.C.M. Kallenberg (eds.), CWI Tract 122, Mathematical Centre, Amsterdam, 1997, pp. 135-149. W.K. Klein Haneveld, O.J. Vrieze and L.C.M. Kallenberg (editors), Ten years LNMB: PH.D. research and graduate courses of the Dutch Network of Operations Research, CWI Tract 122 (1997). G.M. Koole, Stochastic scheduling with event-based dynamic programming. In: 10 years LNMB, W.K. Klein Haneveld, O.J. Vrieze, L.C.M. Kallenberg (eds.), CWI Tract 122, pages 161–169, CWI, Amsterdam, 1997.

Workshops, Seminars a.o.

This chapter summarizes the workshops, seminars and others (co-)organised by (researchers of) the Stieltjes Institute. The following data are given:

- title,

- city and date,

- (co-)organiser(s).

Algebra and Geometry

1.2. Algebraic and Analytic Geometry

AGE Conference Leiden, March 20-23, 1997 M. Lübke, J.P. Murre, A.J.M.M. van der Ven

Conference Strings ’97 Amsterdam, June 16 - 20, 1997 R. Dijkgraaf

Symposium Gravity, Black Holes and Strings Amsterdam, June 21, 1997 R. Dijkgraaf

Conference Low-dimensional varieties in positive characteristic Amsterdam, November 24 - 28, 1997 C. Faber, G. van der Geer

1.4. Theoretical Foundations of Computer Sciences

Graph Grammars and Graph Transformations Day Leiden, June 18, 1997

Analysis

2.1. Operator Theory and Applications

Stieltjes Workgroup Sub-Hardy Hilbert Spaces Amsterdam, November 1996 - March 1997 M.A. Kaashoek, J.J.O.O. Wiegerinck

Symposium Functionaalanalyse en Stochastische Analyse Delft, April 18, 1997 Ph. Clément, J.M.A.M. van Neerven, B. de Pagter

Seminar Probabilistic Methods in Functional Analysis Delft, biweekly Ph. Clément, J.M.A.M. van Neerven, B. de Pagter

One Day Symposium Operator Theory Amsterdam, February 6, 1997 and July 7, 1997 M.A. Kaashoek

Workshop on Dynamics of differential equations with delays Leiden, October 13 - 15, 1997 S.M. Verduyn Lunel

Workshop on Operators and Dynamical Systems Leiden, October 23 - 25, 1997 Ph. Clement, M.A. Kaashoek, S.M. Verduyn Lunel

Workshop Operator Theory and Analysis (at the occasion of the sixtieth birthday of M.A. Kaashoek) Amsterdam, November 12-14, 1997 H. Bart, I. Gohberg and A.C.M. Ran

2.2. Lie Groups, Special Functions, Approximation Theory and Complex Analysis

Seminar Analysis on Symmetric Cones Amsterdam, first half 1997, biweekly G. van Dijk, H.T. Koelink, T.H. Koornwinder

Workshop Special Functions and Group Theory Amsterdam, September 4, 1997 H.T. Koelink

Special Functions Day Amsterdam, December 2, 1997 T.H. Koornwinder

1997 Twente Conference on Lie Groups Enschede, December 15-18, 1997 G. van Dijk, T.H. Koornwinder, E.M. Opdam

Seminar Mathematical Structures in Field Theory Amsterdam, 1997, monthly R.H. Dijkgraaf, N.P. Landsman, H.G.J. Pijls

2.3 Non-linear Partial Differential Equations

Fall Programme Dynamical systems and pattern formation Leiden, September 15 - November 15 1997 L.A. Peletier and S.M. Verduyn Lunel

Workshop Multibump solutions Leiden, October 6 - 9, 1997 A. Doelman, L.A. Peletier and R.C.A.M. van der Vorst

Workshop Interfaces and parabolic regularisation Leiden, November 5 - 7, 1997 C.J. van Duijn en J. Hulshof

Workshop Applied Functional Analysis and Partial Differential Equations Delft, March 25 - 26, 1997

2.4. Numerical Analysis

Stieltjes Analysis Colloquium Leiden, February 27, June 5, and November 10, 1997 K.J. in ’t Hout and J. Hulshof

Seminar Numerical Analysis Leiden; June 5, June 12, August 7, 1997 M.N. Spijker and K.J. in ’t Hout

Mini-symposium on Numerical Analysis Wageningen, April 3 - 4, 1997 P.M. de Zeeuw

TASC Symposium 7 Amsterdam, June 20, 1997 J.G. Verwer

TASC Symposium 8 Amsterdam, September 10, 1997 J.G. Verwer

Mini-symposium on Partial Differential Equations Grado, September 15 - 19, 1997 J.G. Verwer

TASC Symposium 9 Amsterdam, December 11, 1997 J.G. Verwer

Eighth International Seminar Numerical Treatment of Differential Equations Alexisbad, September 1-5, 1997 P.J. van der Houwen

Minisymposium on Parallel ODE methods Grado, September 15-19, 1997 P.J. van der Houwen Working Group Large-Scale Computing Amsterdam, 4 meetings in 1997 H.J.J. te Riele

Second Workshop Cooperative Project Reliability and Parallel Computational Efficiency in Scientific Computing St.Petersburg, Russia, June 27, 1997 P.W. Hemker

3rd Workshop Cooperative Project Reliability and Parallel Computational Efficiency in Scientific Computing Amsterdam, November 19, 1997 P.W. Hemker

International workshop on the Numerical Solution of Thin Layer Phenomena Amsterdam, November 20-21, 1997 P.W. Hemker

Stochastics

3.1. Probability

Heavy tailed distributions and extreme values Rotterdam, December 19, 1997 L. de Haan

XVIII Seminar on Stability Problems of Stochastic Models Debrecen - Hajdúszoboszló, Hungary, January 26 - February 1 1997 J. Mijnheer

3.2. Statistics

FEBS Workshop Analysis of single channel data/Stochastic channel behaviour Leiden, May 3-10, 1997 M.C.M. de Gunst

Semiparametrics and Time Series Leiden, June 30 - July 8, 1997 C.A.J. Klaassen

Operations Research

4.1. Deterministic Optimization

INFORMS annual meeting: Special sections on interior point methods. Dallas, TX, U.S.A., 2-5 May 1997 C. Roos and T. Terlaky

INFORMS spring meeting: Special sections on interior point methods. San Diego, U.S.A., 4-7 May 1997 C. Roos and T. Terlaky

International Workshop High Performance Optimization Techniques II Rotterdam, August 20 - 22, 1997 E. de Klerk, C. Roos, A.J. Quist, T. Terlaky

16th International Symposium on Mathematical Programming: Special sections on interior point methods and on binary matroids. Lausanne, Switzerland, August 24 - 29, 1997 A.M.H. Gerards, M. Laurent, C. Roos and T. Terlaky

4.2. Stochastic Operations Research

Seminar Queueing Theory Amsterdam, June 2 & December 15, 1997 O.J. Boxma

Ninth Summer Workshop of the LNMB Leiden, June 17- 20,1997 L.C.M. Kallenberg Twenty-Second Conference on the Mathematics of Operations Research & Eighth International Workshop of the LNMB Lunteren, January 14 - 16, 1997 A.H.M. Gerards, L.C.M. Kallenberg 93

Foreign Visitors

The following data are given

- name, place and country of the visitor

- name(s) of the host(s)

Algebra and Geometry

1.1. Number Theory

M. Bennett, Ann Arbor, MI, U.S.A., R. Tijdeman A. Beresz, Debrecen, Hungary, R. Tijdeman R. M. Burton, Corvallis, OR, U.S.A., C. Kraaikamp M. van Frankenhuysen, Paris, France, R. Tijdeman L. Hajdu, Debrecen, Hungary, R. Tijdeman T. Langtry, Sydney, Australia, H.J.J. te Riele P. Moree, Bonn, Germany, P. Stevenhagen A.D. Pollington, Provo, UT, U.S.A., R. Tijdeman J.W. Sander, Hannover, Germany, R. Tijdeman Th. A. Schmidt, Corvallis, OR, U.S.A., C. Kraaikamp J. Urbanowicz, Warsaw, Poland, R. Tijdeman

1.2. Algebraic and Analytic Geometry

V. Alexeev, Athens, GA, U.S.A., G. van der Geer A. Besser, Bonn, Germany, J.P. Murre C. Ciliberto, Rome, Italy, G. van der Geer A. Conte, Torino, Italy, J.P. Murre M. Cornalba, Pavia, Italy, G. van der Geer E. Friedlander, Evanston, U.S.A., J.P. Murre K. Galicki, Bonn, Germany, J. Bochnak B. Gordon, Norman, U.S.A., J.P. Murre T. Katsura, Tokyo, Japan, G. van der Geer E. Kucharz, Albuquerque, NM, U.S.A., J. Bochnak A. Moriwaki, Paris, France, G. van der Geer S. Müller-Stach, Essen, Germany, J.P. Murre J. Piontkowski, Düsseldorf, Germany, A.J.H.M. van de Ven

1.3. Topology and Dynamical Systems

J. J. Dijkstra, Tuscaloosa, AL, U.S.A.; J. van Mill M. Husek, Prague, Czech Republic, J. van Mill D. Repovs, Ljubljana, Slovenia, J. van Mill P. Simon, Prague, Czech Republic, J. van Mill

1.4. Theoretical Foundations of Computer Sciences

E. Csuhaj-Varju, Budapest, Hungary, G. Rozenberg T. Harju, Turku, Finland, G. Rozenberg D. Lapoire, Bordeaux, France, J. Engelfriet S. Maneth, Ulm, Germany, J. Engelfriet A. Mateescu, Bucharest, Romania, G. Rozenberg R. Morin, Paris, France, H.C.M. Kleijn G. Paun, Bucharest, Romania, G. Rozenberg L. Rosaz, Paris, France, H.C.M. Kleijn A. Salomaa, Turku, Finland, G. Rozenberg

Analysis

2.1. Operator Theory and Applications

D.Z. Arov, Odessa, Ukraine, M.A. Kaashoek J.-B. Baillon, Lyon, France, Ph.P.J.E. Clément V. Derkach, Donetsk, Ukraine, M.A. Kaashoek Y. Eidelman, Tel-Aviv, Israel, M.A. Kaashoek X. Fernique, Strasbourg, France, Ph.P.J.E. Clément, J. van Neerven A. Frazho, West Lafayette, IN, U.S.A., M.A. Kaashoek J.J. Grobler, Potchefstroom, South Africa, Ph.P.J.E. Clément, B. de Pagter N. Krupnik, Israel, M.A. Kaashoek D. Lamberton, France, Ph.P.J.E. Clément, J. van Neerven, B. de Pagter L. Lerer, Technion, Haifa, Israel, A.C.M. Ran S.-O. Londen, Helsinki, Finland, Ph.P.J.E. Clément J. Mallet Paret, Providence, U.S.A., S.M. Verduyn Lunel C.V.M. van der Mee, Cagliari, Italy, A.C.M. Ran W. Ricker, Australia, B. de Pagter M. Röckner, Bielefeld, Germany, Ph.P.J.E. Clément, B. de Pagter L. Rodman, Williamsburg, U.S.A., A.C.M. Ran O. Staffans, Helsinki, Finland, S.M. Verduyn Lunel J. Sola-Morales, Barcelona, Spain, S.M. Verduyn Lunel F.A. Sukochev, Flinders, Australia, B. de Pagter D. Yakubovich, St-Petersburg, Russia, S.M. Verduyn Lunel J. Zabczyk, Warsaw, Poland, Ph.P.J.E. Clément, J. van Neerven Y. Zucker, Tel-Aviv, Israel, M.A. Kaashoek

2.2. Lie Groups, Special Functions, Approximation Theory and Complex Analysis

R. Alvarez-Nodarse, Madrid, Spain, T.H. Koornwinder N. Ciccoli, Bologna, Italy, T.H. Koornwinder G.L. Litvinov, Moscow, Russia, G. van Dijk V.F. Molchanov, Tambov, Russia, G. Van Dijk M.A. Piñar, Granada, Spain, H.G. Meijer M. Rösler, München, Germany, T.H. Koornwinder H. Rosengren, Lund, Sweden, H.T. Koelink A. Ruffing, München, Germany, T.H. Koornwinder J. Van der Jeugt, Gent, Belgium, H.T. Koelink S.B. Yakubovich, Leuven, Belgium, T.H. Koornwinder M.J. Zygmunt, Warsaw, Poland, T.H. Koornwinder

2.3. Non-lineair Partial Differential Equations

H.W. Alt, Bonn, Germany, J. Hulshof, C.J. van Duijn I. Athanasopoulos, Greece, C.J. van Duijn G.I. Barenblatt, Berkeley, CA, U.S.A., J. Hulshof H. Berestycki, Paris, France, J. Hulshof F. Bernis, Madrid, Spain, J. Hulshof M. Bertsch, Rome, Italy, J. Hulshof, C.J. van Duijn A. Bossavit, France, C.J. van Duijn C.M. Brauner, Bordeaux, France, J. Hulshof C.J. Budd, Bath, U.K., J. Hulshof, C.J. van Duijn B. Buffoni, Bath, U.K., L.A. Peletier, R.C.A.M. van der Vorst R. Camassa, Los Alamos, U.S.A., L.A. Peletier, R.C.A.M. van der Vorst J. Chadam, Pittsburgh, PA, U.S.A., J. Hulshof A.R. Champneys, Bath, U.K., L.A. Peletier, R.C.A.M. van der Vorst G. Dagan, Tel Aviv, Israel, C.J. van Duijn R. DalPasso, Rome, Italy, J. Hulshof, C.J. van Duijn C.N. Dawson, Austin, TX, U.S.A., C.J. van Duijn J.I. Diaz, Spain, C.J. van Duijn C.M. Elliot, Sussex, U.K., J. Hulshof V.A. Galaktionov, Bath, U.K., J. Hulshof M. Garcia-Huidobro, Santiago, Chile, Ph.P.J.E. Clément, G. Sweers H. Garcke, Bonn, Germany, J. Hulshof R.A. Gardner, Amherst, U.S.A., L.A. Peletier, R.C.A.M. van der Vorst J.P. Gossez, Bruxelles, Belgium, Ph.P.J.E. Clément, G. Sweers G. Grun, Bonn, Germany, J. Hulshof B. Grundy, St. Andrews, Scotland, C.J. van Duijn J.K. Hale, Atlanta, GA, U.S.A., L.A. Peletier, R.C.A.M. van der Vorst S.P. Hastings, Pittsburgh, PA, U.S.A., L.A. Peletier, R.C.A.M. van der Vorst G. Heat, CSIRO, Canberra, Australia, C.J. van Duijn D. Hilhorst, Paris, France, L.A. Peletier, R.C.A.M. van der Vorst, J. Hulshof Ph. Holmes, Princeton, U.S.A., L.A. Peletier, R.C.A.M. van der Vorst G.M. Homsy, Stanford, U.S.A., C.J. van Duijn Z. Hongfei, Indiana, U.S.A., C.J. van Duijn B. Kalies, California, U.S.A., L.A. Peletier, R.C.A.M. van der Vorst S. Kamin, Tel Aviv, Israel, J. Hulshof H.G. Kaper, Argonne, U.S.A., L.A. Peletier, R.C.A.M. van der Vorst T. Kaper, Boston, MA, U.S.A., L.A. Peletier, R.C.A.M. van der Vorst T. Kapitula, New Mexico, U.S.A., L.A. Peletier, R.C.A.M. van der Vorst R. Kersner, Paris, France, C.J. van Duijn J.R. King, Nottingham, U.K., J. Hulshof E. Logak, Cergy-Pont, France, L.A. Peletier, R.C.A.M. van der Vorst J. Mallet-Paret, Providence, RI, U.S.A., L.A. Peletier R. Manásevich, Santiago, Chile, Ph.P.J.E. Clément, G. Sweers H. Matano, Tokyo, Japan, L.A. Peletier, R.C.A.M. van der Vorst A. Mielke, Hannover, Germany, L.A. Peletier, R.C.A.M. van der Vorst E. Mitidieri, Trieste, Italy, Ph.P.J.E. Clément, G. Sweers M. Niezgodka, Poland, C.J. van Duijn A. Novick-Cohen, Haifa, Israel, J. Hulshof F. Otto, Bonn, Germany, J. Hulshof M.A. Peletier, Bath, U.K., L.A. Peletier, R.C.A.M. van der Vorst J. Philip, Canberra, Australia, C.J. van Duijn P. Pucci, Perugia, Italy, Ph.P.J.E. Clément, G. Sweers J.F. Rodrigues, Lisbon, Portugal, J. Hulshof, C.J. van Duijn J.-M. Roquejoffre, Toulouse, France, J. Hulshof J. Rubin, Columbus, OH, U.S.A., L.A. Peletier, R.C.A.M. van der Vorst R. Schätzle, Bonn, Germany, L.A. Peletier, R.C.A.M. van der Vorst C. Schmidt-Laine, Lyon, France, J. Hulshof E. Séré, Paris, France, L.A. Peletier, R.C.A.M. van der Vorst J. Sola-Morales, Barcelona, Spain, L.A. Peletier, R.C.A.M. van der Vorst C.A. Stuart, Lausanne, Switzerland, Ph.P.J.E. Clément, G. Sweers D. Terman, Columbus, OH, U.S.A., L.A. Peletier, R.C.A.M. van der Vorst J.F. Toland, Bath, L.A. Peletier, R.C.A.M. van der Vorst J. Trangenstein, Duke University, U.S.A., C.J. van Duijn W.C. Troy, Pittsburgh, L.A. Peletier, R.C.A.M. van der Vorst M. Ughi, Trieste, Italy, J. Hulshof, C.J. van Duijn A. Yew, Providence, RI, U.S.A., L.A. Peletier, R.C.A.M. van der Vorst J.-L. Vazquez, Madrid, Spain, J. Hulshof C. Werner, France, C.J. van Duijn M. Willem, Louvain, Belgium, Ph.P.J.E. Clément, G. Sweers A. de Wit, Bruxelles, Belgium, C.J. van Duijn H.-S. Zhou, Lausanne, Switzerland, Ph.P.J.E. Clément, G. Sweers

2.4. Numerical Analysis

P. Arbenz, Zürich, Switzerland, H.J.J. te Riele R. Baraglia, Pisa, Italy, H.J.J. te Riele U. Becker-Lemgau, GMD-SCAI, Germany, B. Koren C.J. Budd, Bath, U.K., J.G. Verwer G. Carmichael, Iowa, U.S.A., J.G. Verwer M.M. Chawla, Kuwait, P.J. van der Houwen D. Drissi, Kuwait, M.N. Spijker C. Eichler-Liebenow, Halle, Germany, P.J. van der Houwen H. Elbern, Cologne, Germany, J.G. Verwer H.-G. Galbas, Sankt Augustin, Germany, H.J.J. te Riele V. Gerdt, Dubna, Russia, J.A. Sanders C. Greenough, Chilton, U.K., H.J.J. te Riele C. Grossmann, Kuwait, P.W. Hemker R. Hess, R. Lorentz, GMD-SCAI, Germany, B. Koren D. Laforenza, Pisa, Italy, H.J.J. te Riele C.A. Marchi, Rome, Italy, H.J.J. te Riele E. Messina, Napels, Italy, P.J. van der Houwen C. Oosterlee, SCAI, GMD, Germany, B. Koren A. Ostermann, Innsbruck, Austria., K.J. in ’t Hout, M.N. Spijker B. Owren, Trondheim, Norway, K.J. in ’t Hout, M.N. Spijker B. Philippe, Rennes, France, H.J.J. te Riele F. Potra, Iowa, U.S.A., J.G. Verwer K.S. Rao, Madras, India, P.J. van der Houwen A. Schüller, SCAI, GMD, Germany, B. Koren G.I. Shishkin, Ekatierinburg, Russia, P.W. Hemker L.P. Shishkina, Ekatierinburg, Russia, P.W. Hemker G. Söderlind, Lund, Sweden, P.J. van der Houwen U. Trottenberg, SCAI, GMD, Germany, B. Koren B.D. Welfert, Tempe, AZ, U.S.A.., M.N. Spijker

Stochastics

3.1. Probability

M. Aizenman, Princeton, U.S.A., M.S. Keane, J.L. Mijnheer Dr. M. Bloznelis, Vilnius, Lithuania, J.L. Mijnheer L. Bondesson, Uppsala, Sweden, F.W. Steutel T. Bountis, Patras, M.S. Keane A. Bovier, Berlin, Germany, M.S. Keane R. Burton, Corvallis, U.S.A., M.S. Keane S. Cheng, Beijing, China, L. de Haan M. Crowder, Surrey, U.K., T. Bedford P. Deheuvels, Paris, France, L. de Haan H. Drees, Cologne, Germany, L. de Haan H. Farkas, Budapest, Hungary, M.S. Keane Y. Fisher, San Diego, U.S.A., F.M. Dekking, M.S. Keane S. Foss, Novosibirsk, Russia, G. Hooghiemstra I. Fraga Alves, Lisbon, Portugal, L. de Haan S. French, Leeds, U.K., T. Bedford J.-F. le Gall, Paris, France, M.S. Keane D. Gilat, Tel Aviv, Israel, P.J. Holewijn, Israel, M.S. Keane C.M. Goldie, Sussex, U.K., F.W. Steutel A. Greven, Erlangen, Germany, M.S. Keane G. Grimmett, Cambridge, U.K., M.S. Keane T.P. Hill, Atlanta, GA, U.S.A., P.J. Holewijn, M.S. Keane M. Hollander, Seattle, U.S.A., F.M. Dekking, C. Kraaikamp A.-J. Homburg, Berlin, Germany, M.S. Keane M. Huśková, Prague, Czech Republic, M.S. Keane J. Hüssler, Bern, Switzerland, G. Hooghiemstra A. Iwanik, Wroclaw, Poland, M.S. Keane S. Karia, Pune, India, T.J. Bedford V. Kalashnikov, Moscow Russia G. Hooghiemstra O.I. Klesov, Kiev, Ukrain, J.L. Mijnheer H. Kuiper, Tuscon, AZ, U.S.A., M.S. Keane Y. Lacroix, Brest, France, M.S. Keane G. Last, Bonn, Germany; G. Hooghiemstra J. Levy-Vehel, Roquencourt, France, F.M. Dekking M. Lifshits, St. Petersburg, Russia, J.L. Mijnheer V. Lotov, Novosibirsk, Russia, M.S. Keane K. Marton, Budapest, Hungary, J.L. Mijnheer, M.S. Keane J. Mémin, Rennes, France, M.S. Keane M. Menshikov, Moscow, Russia, M.S. Keane P.A. Mykland, Chicago, U.S.A., L. Meester A. Nobel, Chapel Hill, U.S.A., M.S. Keane H. Osinga, Minneapolis, U.S.A., M.S. Keane E.I. Pancheva, Sofia, Bulgaria, A.A. Balkema T.Th. Perreira, Lisbon, Portugal, L. de Haan V. Piterbarg, Moscow, Russia, L. de Haan Y. Qi, Peking, P.R. China, L. de Haan R. Reiss, Siegen, Germany, L. de Haan S. Resnick, Ithaca, U.S.A., L. de Haan H. Rootzén, Göteborg, Sweden, L. de Haan R. Roy, New Delhi, India, M.S. Keane S. Schmid, Berlin, Germany, T. Bedford T.A. Schmidt, Corvallis, OR, U.S.A., M. Keane, C. Kraaikamp V. Schmidt, Ulm, Germany G. Hooghiemstra J. Serafin, Wroclaw, Poland, M.S. Keane V. Sirvent, Caracas, Venezuela, F.M. Dekking J.Q. Smith, Warwick, U.K., T. Bedford M. Smorodinsky, Tel Aviv, Israel, M.S. Keane J. Tawn, Lancaster, U.K., A.K. Sinha B. Tóth, Budapest, M.S. Keane K.F. Turkman, Lisbon, Portugal, L. de Haan V.Vatutin, Moscow, Russia, F.M.Dekking H. Xin, Beijing, China, L. de Haan R. Zitikis, Ottawa, Canada, M.S. Keane

3.2. Statistics

P.J. Bickel, Berkeley, U.S.A.. W.R. Van Zwet G.R. Grimmett, Cambridge, U.K. W.R. Van Zwet M. Huskova, Prague, Czech Republic, W.R. Van Zwet H.-R. Künsch, Zürich, Switzerland, M.C.M. de Gunst, A.W. van der Vaart D.E. Loeb, Bordeaux, France, A. Di Bucchianico E.G. Luebeck, Seattle, WA, U.S.A., M.C.M. de Gunst, A.W. van der Vaart S. Murphy, Pittsburgh, PA, U.S.A., A.W. van der Vaart A. Nobel, Chapel Hill, NC, U.S.A., A.W. van der Vaart J. Oeppen, Cambridge, U.K., A.W. van der Vaart J. Palmgren, Stockholm, Sweden, M.C.M. de Gunst, A.W. van der Vaart J. Robins, Boston, MA, U.S.A., A,W, van der Vaart G.-C. Rota, Cambridge, MA, U.S.A., A. Di Bucchianico F.H. Ruymgaart, Lubbock, TX, U.S.A.., J. Einmahl J.A. Wellner, Seattle, WA, U.S.A., A.W. van der Vaart R.Zitikis, Ottawa, Canada, R. Helmers

Operations Research

4.1. Deterministic Optimization

E.D. Andersen, Odense, Denmark, C. Roos, T. Terlaky K. Anstreicher, Louvain-la-Neuve, Belgium, C. Roos, T. Terlaky I. Costello, Costa Rica, C. Roos R. Freund, Cambridge, Mass., U.S.A., C. Roos, T. Terlaky K. Fukuda, Zürich, Switzerland, T. Terlaky C.C. Gonzaga, Florianopolis, Brazil, C. Roos, T. Terlaky T. Illés, Budapest, Hungary, C. Roos, T. Terlaky A.V. Karzanov, Moscow, Russia, A. Schrijver G. Kassay, Cluj, Romania, T. Terlaky F.R. McMorris, Louisville, KT, U.S.A., H.M. Mulder J. Mitchell, Troy, NY, U.S.A., C. Roos, T. Terlaky H. Wolkowicz, Waterloo, Canada, C. Roos, T. Terlaky

4.2. Stochastic Operations Research

E. Altman, Sophia Antipolis, France, A. Hordijk V. Feltkamp, Alicante, Spain, A. Hordijk, L.C.M. Kallenberg B. Gaujal, Sophia Antipolis, France, A. Hordijk V.A. Malyshev, INRIA, Rocquencourt, Frankrijk, A. Hordijk, F.M. Spieksma Ph. Nain, INRIA, France, O.J. Boxma D. Perry, Haifa, Israel, O.J. Boxma M.L. Puterman, British Columbia, Canada, A. Hordijk, L.C.M. Kallenberg V. Schmidt, Ulm, Germany, O.J. Boxma A. Shwartz, Haifa, Israel, A. Hordijk K. Sigman, U.S.A., O.J. Boxma L. Thomas, Edinburgh, U.K., A. Hordijk, L.C.M. Kallenberg A.A. Yushkevich, North Carolina, U.S.A. A. Hordijk G. Weiss, Haifa, Israel, O.J. Boxma 103

Research Staff

(Situation at July 1, 1997)

Algebra and Geometry

1.1. Number Theory

Permanent staff

Dr. J. Brinkhuis EUR Dr. J.-H. Evertse RUL Dr. C. Kraaikamp TUD Dr. P. Stevenhagen UvA Dr. R.J. Stroeker EUR Prof.dr. R. Tijdeman RUL Dr. R.W. van der Waall UvA

Post Doc’s

Dr. B. de Smit UvA - NWO Dr. B.M.M. de Weger RUL - NWO

Ph.D. students

drs. N.R. Bruin RUL drs. S. Cavallar RUL/CWI-NWO drs. A.C.P. Gee UvA drs. A. Heinis RUL drs. H. Roskam UvA - NWO

CWI participants

Dr.ir. H.J.J. te Riele

1.2. Algebraic and Analytic Geometry

Permanent staff

Prof.dr. J. Bochnak VU Prof.dr. R.H. Dijkgraaf UvA Prof.dr. G.B.M. van der Geer UvA Dr. M. Lübke RUL drs. M.A.J.G. van der Vlugt RUL

Post Doc’s

Dr. N.P. Landsman UvA - KNAW Dr. J.H. Piontkowski RUL Dr. M. Serone UvA - NWO

Ph.D. students

drs. M. Boguslavski UvA drs.J.-H. Cremers UvA - NWO drs. M. Dekker UvA drs. G. Farkas UvA - NWO drs. J. van Hamel VU C.D. Neumann UvA drs. V. Shabat UvA - NWO drs. C.G. Zaal UvA

1.3. Topology and Dynamical Systems

Permanent staff

Prof.dr. J.M. Aarts TUD Dr. A.J.M. van Engelen EUR Dr. K.P. Hart TUD Prof.dr. J. van Mill VU Prof.dr. S.J. van Strien UvA Dr. M.L.J. van de Vel VU

Ph.D. students

drs. H.S. Brandsma VU ir. M.A. van Hartskamp VU O. Kozlovsky UvA

1.4. Theoretical Foundations of Computer Sciences

Permanent staff

Dr. J. Engelfriet RUL Dr. H.J. Hoogeboom RUL Dr. H.C.M. Kleijn RUL Prof.dr. G. Rozenberg RUL

Ph.D. Students

drs. M. ter Beek RUL drs. T. Gelsema RUL drs. J. Hage RUL drs. R. van Vliet RUL drs. N. van Vugt RUL

CWI participants

Prof.dr. J.W. de Bakker Dr. J.J.M.M. Rutten

2.1. Operator Theory and Applications

Permanent staff

Prof.dr. H. Bart EUR Prof.dr. Ph.P.J.E. Clement TUD Prof.dr. I. Gohberg VU Prof.dr. M.A. Kaashoek VU Dr. B. de Pagter TUD Dr. A.C.M. Ran VU Dr. F. van Schagen VU Dr. G.Ph.A. Thijsse EUR Prof.dr. S.M. Verduyn Lunel UvA - KNAW

Post Doc’s

Dr. J. van Neerven TUD - KNAW

Ph.D. students

drs. B. Lemmens UvA/CWI - NWO drs. D.R. Pik VU

2.2. Lie Groups, Special Functions, Approximation Theory and Complex Analysis

Permanent staff

Dr. H. Bavinck TUD Dr. M.G. de Bruin TUD Prof.dr. G. van Dijk RUL Dr. E. Hendriksen UvA Dr. R. Koekoek TUD Prof.dr. T.H. Koornwinder UvA Prof.dr. H.G. Meijer TUD Dr. E.M. Opdam RUL Dr. P.J.I.M. de Paepe UvA Dr. H.G.J. Pijls UvA Dr. J.J.O.O. Wiegerinck UvA

Post Doc’s

Dr. H.T. Koelink UvA - NWO Dr. Liu Heping UvA Dr. A. Pasquale RUL - NWO

Ph.D. students

drs. P.R. Beneker UvA drs. S.C. Hille RUL ir. P.J. Oonincx UvA/CWI - NWO drs. Y.A. Sharshov RUL - NWO drs. J. Stokman UvA

CWI participants

Dr. N.M. Temme Dr. R.A. Zuidwijk STW

2.3. Non-lineair Partial Differential Equations

Permanent staff

Prof.dr. Ph.P.J.E. Clement TUD Prof.dr. C.J. van Duijn TUD/CWI Dr. J. Hulshof RUL Dr. J.F. Kaashoek EUR Prof.dr.ir. L.A. Peletier RUL Dr. G.H. Sweers TUD

Post Doc’s

Dr. B.W. van der Fliert RUL - NWO

Ph.D. students

drs. J.B. van den Berg RUL drs. A.I. Rotariu - Bruma RUL - NWO ir. R.J. Schotting TUD/CWI ir. M.F. Uiterdijk TUD

2.4. Numerical Analysis

Permanent staff

Prof.dr. P.W. Hemker UvA/CWI Dr.ir. G.C. Herman TUD Prof.dr. P.J. van der Houwen UvA/CWI Dr. P.R. Pfluger UvA Dr. J.A. Sanders VU Prof.dr. M.N. Spijker RUL Prof.dr. M. van Veldhuizen VU

Post Doc’s

Dr. K.J. in ’t Hout RUL - KNAW Dr. S. Tracogna RUL

Ph.D. students

drs. N. Borovykh RUL drs. E.G. van den Heuvel RUL ir. J. Noordmans UvA/CWI drs. E.J. Spee UvA/CWI drs. J.J.B. de Swart UvA/CWI - NWO drs. J. Wang VU

CWI participants

Dr. W.H. Hundsdorfer Dr.ir. B. Koren Dr.ir. H.J.J. te Riele Dr. B.P. Sommeijer Dr. J.G. Verwer

Stochastics

3.1. Probability

Permanent staff

Dr. A.A. Balkema UvA Dr. T. Bedford TUD Prof.dr. R.M. Cooke TUD Prof.dr. F.M. Dekking TUD Dr. J.L. Geluk EUR Prof.dr. L.F.M. de Haan EUR Dr. K. van Harn VU Dr. G. Hooghiemstra TUD Prof.dr. M.S. Keane TUD/UvA/CWI Dr.ir. M. Keijzer TUD Dr. C. Kraaikamp TUD Dr.ir. L.E. Meester TUD Dr. J.L. Mijnheer RUL

Ph.D. students

drs. P.C. Allaart VU drs. Ashoke Ermakov TUD/CWI - NWO drs. Liang Peng EUR drs. B.A.M. Schouten UvA/CWI drs. A. Sinha EUR - NWO

CWI participants

Dr. J van den Berg Dr. M.N.M. van Lieshout

3.2. Statistics

Permanent staff

Dr.ir. E. van Berkum TUE Dr. A. di Bucchianico TUE Prof.dr. R.J.M.M. Does UvA Dr. J. Einmahl TUE Dr. A.J. van Es UvA Dr. S.A. van de Geer RUL Prof.dr. P. Groeneboom TUD Dr. M.C.M. de Gunst VU Prof.dr. L.F.M. de Haan EUR Dr.ir. G. Jongbloed TUD Prof.dr. C.A.J. Klaassen UvA Dr. A.J. Koning EUR Prof.dr. P. van der Laan TUE Dr. H.P. Lopuhaä TUD Dr. A.W. van der Vaart VU Prof.dr. W.R. van Zwet RUL

Post Doc’s

drs. S.A. Vogelzang VU

Ph.D. students

drs. A.F. Huele UvA drts. R.A. Ion UvA - NWO drs. M.A. Jonker VU drs. A.J. Lenstra UvA drs. J.J. Lok VU drs. I.W. Mangka TUD/CWI drs. N. Mushkudiani TUE ir. J.G. Schouten VU ir. M.A. van de Wiel TUE

CWI participants

Dr. K. Dzhaparidze Dr.ir. H.J.A.M. Heijmans Dr. R. Helmers

Operations Research

4.1. Deterministic Optimization

Permanent staff

Prof.dr. A. Blokhuis VU Prof.dr. F.A. Lootsma TUD Dr. H.M. Mulder EUR Dr.ir. C. Roos TUD Prof.dr. A. Schrijver UvA/CWI Dr. T. Terlaky TUD

Ph.D. students

drs. T. Fleiner UvA/CWI - NWO drs. J. Keijsper UvA ir. E. de Klerk TUD drs. R. Pendavingh UvA ir. A.J. Quist TUD ir. H. Schuijt TUD

CWI participants

Dr.ir. A.M.H. Gerards

Permanent staff

Dr.ir. I.J.B.F. Adan TUE Prof.dr. R.M. Cooke TUD Prof.dr. A. Hordijk RUL Prof.dr. L.C.M. Kallenberg RUL Dr. G.M. Koole VU Dr. J. Resing TUE Dr.ir. J. van der Wal TUE Prof.dr. J. Wessels TUE

Post Doc’s

Dr. F.M. Spieksma RUL - KNAW

Ph.D. students

ir. B.C.P. Kraan TUD drs. N. Popov RUL

CWI participants

Prof.dr. O.J. Boxma Prof.dr. F.A. van der Duyn Schouten

Organization of the Thomas Stieltjes Institute

(Situation at July 1, 1997)

Board

Prof.dr. M.A. Kaashoek (VU), chairman Prof.dr. J.M. Aarts (TUD) Prof.dr. H. Bart (EUR) Prof.dr. G. van Dijk (RUL) Prof.dr. G.B.M. van der Geer (UvA) Dr. G. Zoutendijk Dr. N.M. Temme (CWI), advisor Prof.dr. J. Wessels (TUE), advisor

Managing board

Prof.dr. W.R. van Zwet (RUL), scientific director Drs. F. Bakker (RUL), secretary

Science Committee

Board: Prof.dr. R. Tijdeman (RUL), chairman Dr. F.M. Dekking (TUD) Prof.dr. T.H. Koornwinder (UvA) Prof.dr. J. Wessels (TUE) Members: Prof.dr. H. Bart (EUR) Prof.dr. P.J. van der Houwen (UvA/CWI) Prof.dr. J. van Mill (VU) Prof.dr.ir. L.A. Peletier (RUL) Prof.dr. A. Schrijver (UvA/CWI) Prof.dr. A.W. van der Vaart (VU)

Prof.dr. A.J.H.M. van de Ven (RUL)

Education Committee

Board: Prof.dr. J.M. Aarts (TUD), chairman Dr. M.G. de Bruin (TUD), secretary/coordinator Members: Prof.dr. T.H. Koornwinder (UvA) Prof.dr. A. Schrijver (UvA/CWI) Prof.dr. M.N. Spijker (RUL) Dr. A.W. van der Vaart (VU)